### v1

# RT-GuIDE: Real-Time Gaussian splatting for Information-Driven Exploration

Abstract
We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing information-rich maps. Further, we develop a parallelized motion planning algorithm that can exploit the Gaussian map for real-time navigation. The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy. We show through simulation experiments that our method is competitive with approaches that use alternate information gain metrics, while being orders of magnitude faster to compute. In real-world experiments, our algorithm achieves better map quality (10% higher Peak Signal-to-Noise Ratio (PSNR) and 30% higher geometric reconstruction accuracy) than Gaussian maps constructed by traditional exploration baselines. Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
I Introduction
Active mapping is a problem of optimizing the trajectory of an autonomous robot in an unknown environment to construct an informative map in real-time. It is a critical component of numerous real-world applications such as precision agriculture [1], infrastructure inspection [2], and search and rescue [3] missions. While nearly all tasks rely on recovering accurate metric information to enable path planning, many also require more fine-grained information. Recent advances in learned map representations from the computer vision and graphics communities [4, 5] have opened up new possibilities for active mapping and exploration while maintaining both geometrically and visually accurate digital twins of the environment.
While prior work effectively solves the problem of information-driven exploration [6, 7, 8, 9, 10] or frontier-based exploration [11, 12, 13], in this work we consider the additional problem of generating radiance fields while also performing autonomous navigation. Prior work has also proposed information metrics using novel learned scene representations that are capable of high quality visual reconstruction but are incapable of running in real-time onboard a robot. To enable efficient mapping and planning in these novel representations, we consider approximation techniques for computing the information gain. Further, we consider the problem of generating high-quality maps that are capable of novel-view synthesis that can be used for downstream applications.
Fig. 1 shows the elements of our approach. We use the Gaussian splatting approach proposed in [14] to generate Gaussian maps. We propose a novel information gain metric to compute the utility of regions in the environment and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path. Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
Figure 1: Key elements of our proposed approach. [A] Robot building a Gaussian map onboard in real-time and using it to avoid obstacles in the environment. Synthesized color and depth images from the Gaussian map are presented next to the corresponding observations from the RGBD sensor. [B] Robot navigating to unobserved areas (right) with high information gain while maximizing information along the trajectory. Gaussian frontiers colored cyan (light to dark) in increasing estimated information gain.
In summary, the contributions of this paper are:
  1. 1.
    A unified framework for online mapping and planning with Gaussians.
  2. 2.
    An approximate information gain metric that is easy to compute and capable of running in real-time onboard an unmanned ground vehicle.
  3. 3.
    A method to compute frontiers in a map built using Gaussian splatting.
  4. 4.
    Extensive experiments in both indoor and outdoor environments to evaluate our framework.
II Related Work
Map Representation. To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community. The most intuitive but effective volumetric representation has been widely used. Voxel-based representation could maintain information such as occupancy [15] or signed distance [16, 17]. With the recent application of semantic segmentation, semantic maps that contain actionable information have been proposed [18, 19, 20, 21, 22]. With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [4] and 3D Gaussian Splatting (3DGS) [5] have become popular representations for robotic motion planning. In this work, we study the problem of active mapping with the learned map representations.
Active Simultaneous Localization and Mapping (SLAM). The problem of exploration and active SLAM has been widely studied in the past decade. The classical exploration framework uses a model-based approach to actively navigate towards frontiers [11, 12] or waypoints that have the highest information gain [6, 7, 8, 9, 10, 23, 24]. Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [25, 26, 27, 28]. However, most of the existing work developed their approaches based on classical map representations. For example, frontiers are typically defined in voxel maps, information metrics are typically associated with occupancy maps or Signed Distance Field (SDF) maps.
In this work, we instead consider the active mapping problem with a learned map representation. Bayesian neural networks [29, 30] and deep ensembles [24, 31] are common approaches for estimating uncertainty in learned representations. [32, 33] use the idea of computing the uncertainty based on consecutive measurements from a moving robot in contrast to train multiple models in the traditional ensemble frameworks. Radiance field representations provide additional possibilities for estimating uncertainty through the volumetric rendering process [34]. [35] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty.  [36] leverages Fisher information to compute pixel-wise uncertainty on rendered images. Where prior work in implicit and radiance field representations necessitates use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a Gaussian map representation encodes physical parameters of the scene, which motivates estimating information gain from the Gaussians directly.
Navigation in Radiance Fields. Prior work has also considered planning directly in radiance fields. [37] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration. Sim-to-real approaches leverage the rendering quality of the learnt radiance fields to train downstream tasks such as visual localization, imitation learning [38] by effectively utilizing the representation as a simulator. [39] use a liquid neural network to train imitation policies for a quadrotor robot in a Gaussian splat. [40] utilize the geometric fidelity of the representation to first map the environment and then perform trajectory optimization to plan paths in these environments. [41] use a pre-generated Gaussian map to compute safe polytopes to generate paths and also present a method to synthesize novel views with a coarse localization of the robot and refine the estimate by solving a Perspective-n-Point (PnP) problem. The probabilistic representations of free space has also been utilized for motion planning [42] where authors use uncertainties in the learned representation to provide probabilistic guarantees on collision-free paths.
III Problem Specification
Figure 2: The proposed active mapping framework. The proposed framework contains two major components, the planning module and the mapping module. As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation $m$ at every step and computes the utility of geometrically clustered frontiers. The information is then passed to the planning module which comprises the topological graph and motion primitive library ([B]). The topological graph adds sampled viewpoints as nodes and passes along a planned path to the trajectory planner. The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper). The planned trajectory is executed by the robot to get a new set of observations.
We have an unknown map of the environment $m^{*}$ that belongs to a space of maps $\mathcal{M}$. Our goal is to solve the following problem
Problem 1.
Information-Driven Exploration.
$\min_{x_{1:T}}$ $\ \mathbb{E}_{x}[\rho(h(x,\hat{m}),h(x,m^{*}))]$ (1)
$s.t$ $\ \hat{m}=\Phi(y_{1:T},x_{1:T})$
$\ y_{t}=h(x_{t},m^{*})\ \forall 1\leq t\leq T.$
In the equation above, the expectation is taken over a random evaluation point $x$. The states belong to the state space $\mathcal{X}$, whereas observations belong to the observation space $\mathcal{Y}$. The observation function $h:\mathcal{X}\times\mathcal{M}\rightarrow\mathcal{Y}$ takes as input the robot state together with the map of the environment, and outputs the rendered image. $T$ represents the exploration budget. $m^{*}$ is the true map of the environment. $\hat{m}$ is the map of the environment that is estimated from the sequence of measurements $(y_{t})_{1\leq t\leq T}$ taken from vantage points (states) $(x_{t})_{1\leq t\leq T}.$ The discrepancy function $\rho$ captures the difference between the image rendered based on the estimated map, and the actual map. The sequence of vantage points is constrained to lie in free space $\mathcal{X}_{free}$.
Since Prob (1) is parameterized by $m^{*}$, which is a priori unknown, the problem is ill-posed. We are instead interested in synthesizing a (near) optimal policy $(\pi_{t})_{t=0}^{T}$, consisting of a sequence of functions that map the history of observations to the optimal action to take at the current step. In particular, for any $i\leq T$, the function
$\pi_{t}:\mathcal{Y}_{1}\times\cdots\times\mathcal{Y}_{t}$ $\rightarrow\mathcal{X}$ (2)
$\pi_{t}(y_{1},y_{2},...,y_{t})$ $=x_{t}$
ought to specify the next vantage point given the the history of observations accrued thus far. Finding the optimal sequence of policies is a challenging, high-dimensional, optimization problem. We therefore develop an approximate method for solving it. The approximation lies in a particular, but suitable, choice of representation for $m^{*}$, together with a way of using it for effective exploration.
IV Method
Our proposed framework comprises the mapping module and the planning module, as illustrated in Fig. 2. The mapping module (Sec. IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the utility of Gaussian frontiers. The information is then passed to the planning module for planning guidance paths (Sec. IV-B1) and trajectories (Sec. IV-B2).
IV-A Mapping & Uncertainty Estimation
IV-A1 Mapping
We adopt the 3D Gaussian Splatting (3DGS) approach to represent the environment as a map of 3D Gaussians. 3DGS enables real-time radiance field rendering by projecting and blending the 3D Gaussians to render images. The parameters of the 3D Gaussians are optimized through mapping iterations to improve the representation of the scene. We build the 3DGS mapping module upon the [14] framework. The poses of the robot are obtained using [43].
IV-A2 Gaussian uncertainty for viewpoint selection
We estimate uncertainty in the map representation directly on the 3D Gaussians. We motivate our metric of information gain leveraging the theory of Kalman filtering. In particular, the updates of the Kalman filter [44] are given by
$m_{k|k}$ $=m_{k|k-1}+K(y_{k}-Hx_{k|k-1})$ (3)
$K$ $=P_{k|k-1}H^{T}S^{-1}$
$S$ $=HP_{k|k-1}H^{T}+R$
$P_{k|k}$ $=P_{k|k-1}-KHP_{k|k-1}.$
where $K$ is the Kalman Gain, $P$ is the covariance of the map, $H$ is the linearized measurement model, $S$ is the innovation covariance and $R$ is the covariance of the measurement noise. The following identity holds:
$Cov_{k|k-1}(m_{k|k}-m_{k|k-1})=P_{k|k-1}-P_{k|k}$ (4)
The pertinent aspect of the relation above is that the prior covariance of the change in the map parameters upon observing a new measurement is equal to the information gain of the latter. Although this argument can be rigorously shown in the context of linear measurement models, we nevertheless use it as a measure of future information gain in the context of measurements with a Gaussian representation of the environment.
Gaussian uncertainty. Following the above motivation, we compute the change in Gaussian parameters across each mapping optimization and use the magnitude of the parameter updates $\|\hat{m}_{t}-\hat{m}_{t-1}\|$ as a measure of uncertainty. Gaussians with large changes in parameters represent parts of the scene with high uncertainty and conversely Gaussians with stable parameters have low uncertainty. We use this to approximate information gain for viewpoint selection.
Given a map $m$ represented by a set of Gaussians $G$, we classify the Gaussians in the scene $G=\{G_{H},G_{L},G_{O}\}$ into three disjoint subsets, $G_{H}$ which consists of Gaussians with high uncertainty, $G_{L}$ which consists of Gaussians with low uncertainty and $G_{O}$ for the rest. For a Gaussian $g_{i}\in G$ and pose $x_{t}$ at timestep $t$, we define the binary visibility function $v(x_{t},g_{i})$ capturing whether $g_{i}$ is in the field of view of $x_{t}$.
To maximize information gain while exploring the environment, we aim to find the viewpoint with the highest utility $\xi(x_{t})$ that maximizes the number of high-uncertainty Gaussians and minimizes the number of low-uncertainty Gaussians in the field of view at each timestep. The first term in Equ. 5 encourages observations of parts of the scene with high uncertainty while the second term encourages exploration of unseen areas of the scene. The balance of exploitation and exploration is weighted by the term $\lambda_{\xi}$.
$\xi(x_{t})=\sum_{i\in L}v(x_{t},g_{i})-\lambda_{\xi}\sum_{i\in H}v(x_{t},g_{i})$ (5)
Gaussian frontiers. In contrast to traditional mapping representations, the Gaussian map does not encode occluded and free space. Instead of frontiers of unobserved regions, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next, since Gaussians that have few observations or lie at the edges of observed regions naturally have higher uncertainty. For a region $o\subset m$, we consider the set of Gaussians $\{g\sim\mathcal{N}(\mu,\sigma^{2})\in G\ |\ \mu\in o\}$ with cardinality $N_{g}$ in the region and compute the mean uncertainty of that frontier
$\Omega=\frac{1}{N_{g}}\sum^{N_{g}}_{i}\|\mu^{t}_{i}-\mu^{t-1}_{i}\|.$ (6)
This provides a unified framework for estimating information gain to optimize map quality while implicitly identifying frontiers to guide the exploration of unknown regions.
IV-A3 Traversable region segmentation
We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor. This necessitated the removal of Gaussians on the ground for planning. This also fulfilled an important role in reducing the number of Gaussians that have to be queried in the collision check.
IV-B Hierarchical Planning
We represent the planning problem as a hierarchical planning problem. The first level provides guidance to map particular regions of the environment and a path to that region from the known space in the environment. The second level finds a trajectory that is dynamically-feasible (i.e. obeys the robot’s physical constraints); collision-free and locally maximizes the information along the path.
IV-B1 High-level guidance
We formulate a high-level guidance path that allows us to (i) plan to transition to high uncertainty regions of space which might involve traveling through those with low uncertainty; and (ii) estimate the traversable space in the known space without explicitly having to plan long-range trajectories with computationally expensive collision checks. At the higher level of our planner, we construct a graph $\mathcal{G}(N,E)$ by incrementally adding nodes along the traveled path. The graph consists of two types of nodes $n_{o},n_{p}$ where $n_{o}$ are observed nodes (along our odometry) and $n_{p}$ are predicted nodes. At each planning iteration, we sample a fixed number of viewpoints around the identified regions. We can compute the shortest viewing distance from the Gaussian frontier $f\in\mathbb{R}^{3}$ to the optical center of the camera as
$d_{1}=\frac{||f^{T}\left[b_{1}\ b_{2}\right]||_{2}}{tan\left(\phi/2\right)}$
where $b$ is the camera axis and $\phi$ is the field of view of the camera. Each sampled viewpoint is then assigned the utility computed by the Gaussian frontiers and connected to the closest point in the graph. We then use Dijkstra’s search algorithm to find the shortest path from the current robot location in the graph to all the predicted nodes in the graph. We then compute the cost-benefit of a path using $\Omega/e^{d_{2}}$ where $\Omega$ is the estimated information gain and $d_{2}$ is the distance to the node [45]. The maximal cost-benefit path is sent to the trajectory planner.
IV-B2 Trajectory Generation
The second level planning problem involves finding a collision-free trajectory from start to goal that meets the dynamic constraints of the robot while optimizing the information along the path. The problem is defined as follows:
Problem 2.
Trajectory Planning. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
$\min_{u(\cdot),\tau}\ J(u(\cdot),\tau)$ (7)
$s.t.\ \text{for all }t\in[0,\tau]$
$\dot{x}(t)=f(x(t),u(t)),\ x(t)\in\mathcal{X}_{free},$
$||v(t)||_{2}\leq v_{max},\ |\omega(t)|\leq\omega_{max},$
$x(0),\theta(0)=x_{0},\theta_{0},\ x(\tau)\in\mathcal{X}_{goal}$
where $J$ is a function that encodes a positive combination of execution time $\tau$ and the control effort $J_{c}(u(\cdot))$.
We use the unicycle model as the robot dynamics. The robot state $x=[p,\theta]\in\mathbb{R}^{2}\times S^{1}$ consists of its position ($p$) and heading ($\theta$). The control inputs $u=[v,\omega]\in\mathbb{R}^{2}$ consist of linear velocity ($v$) and angular velocity $(\omega)$. Inspired by [46], we solve problem 2 by performing a tree search on the motion primitives tree. Motion primitives are generated with fixed control inputs over a time interval $\delta\tau$ with known initial states. Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $[0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs. Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization $d\tau$.
We sample a fixed number of points on each motion primitive to conduct collision checks. Since we have an uncertain map, we relax $x\in X_{free}$, to a chance constraint. First, we measure the distance from the discretized positions
$w\sim\mathcal{N}(\kappa,r_{robot}^{2})\in W|\kappa=x_{t}$
along the search path to all Gaussians
$\{g\sim\mathcal{N}(\mu,\sigma^{2})\in G\}$
in the map. Our constraint then boils down to the probability of the distance between the test point and the set of Gaussians in the scene
$d_{g}=\inf||W-G||$
being less than 0 i.e. $P(d_{g}<0)\leq\eta$ where $\eta$ is the tolerable risk. When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $\sigma$ of each Gaussian is scaled by a factor of $\lambda_{g}$. The truth value of the test is determined by comparing the distance to the closest Gaussian with $r_{robot}+\lambda_{g}\sigma$, as illustrated in Fig. 2. The rationale for the latter test is as follows. Assuming that the closest obstacle is unique, the probability of collision with some obstacle is well-approximated by the probability of collision with the nearest obstacle. This holds by virtue of the exponential decay of the Gaussian distribution with the distance from the origin. As a result, specifying an upper bound on the probability of collision is approximately the same as specifying the probability of collision with the closest obstacle.
We develop a GPU-accelerated approach for testing all sampled points with all Gaussians from the map at once while growing the search tree. This allows the real-time expansion of the search tree. The cost of each valid motion primitive is defined as
$J(u(\cdot),\tau):=\lambda_{t}\tau+(v^{2}+w^{2})\tau,$
where $\lambda_{t}$ weights the time cost with the control efforts. Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as
$h(p):=||p_{goal}-p||_{2}/v_{max}.$
We use A* to search through the motion primitives tree and keep top $k$ candidate trajectories for the subsequent information maximization. The information of each trajectory is evaluated with state sequence $[x(\delta\tau),\ldots,x(\tau)]$ on the trajectory according to (5). Finally, the trajectory with the highest information $\left(\sum_{t=0}^{\tau}\xi(x(t))\right)$ is selected and executed by the robot.
IV-C Implementation Details
We implemented the proposed method using PyTorch2.4. Most of our parameters of the Gaussian map were set to the default configuration of [14] for the TUM dataset. For simulation experiments, we used 50 mapping iterations. For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence. The mapping module ran at 1Hz and the planning module ran at 3Hz onboard the robot. We set Gaussian frontier size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$.
V Simulation Results & Ablations
Table I: Simulation Results on 120 iTHOR scenes with 20 steps
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$ t/step $\downarrow$
[dB] [m] [s]
Ensemble 12.589 0.483 0.612 1.064 11.094
FisherRF 13.702 0.558 0.549 1.016 5.347
Ours 13.591 0.533 0.576 0.975 0.046
Environments. We conduct simulation experiments with the AI2-THOR simulator [47] on the iTHOR dataset. The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms.
Evaluation metrics & Baselines. For each scene, we generate a test set of images and evaluate the rendered images from each method against the test set. We evaluate the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [48], Learned Perceptual Image Patch Similarity (LPIPS) [49] on the RGB images and Root Mean-Square-Error (RMSE) on the depth images.
We evaluate the proposed information gain metric against two other view selection methods. For the Ensemble baseline, we train an ensemble of $5$ models with the leave-one-out procedure at each step. We compute the patch-wise variance across the rendered images for each sampled viewpoint and ensemble model, and select the viewpoint with the highest variance as the next action. For FisherRF, we follow the original implementation to evaluate viewpoints based on the Fisher information.
Experiments and results. In our simulation experiments, we set the grid size of reachable positions to 0.25m. At each step, we sample viewpoints with different yaw angles around the current position of the agent as possible actions to take. These viewpoints are then evaluated with the corresponding uncertainty metric and the viewpoint with the highest uncertainty is selected as the next action. We ran experiments on all 120 scenes from iTHOR with 20 steps each. We synthesized novel views according to the test set poses and computed the image metrics against the test set images.
The averaged results are presented in Tab. I. Our information gain metric performs comparably with FisherRF and
Figure 3: Comparison of viewpoint selection methods on environment FloorPlan7 of the iTHOR dataset with 50 steps. The average and variance are computed across 5 different seeds. Our proposed method achieves comparable performance to the other methods while exhibiting significantly smaller variance across runs.
outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time. Since we are computing the uncertainty directly on the Gaussians instead of the image, we avoid the computation cost of rendering every sampled viewpoint. This efficiency in computation is crucial in a sampling-based planner where we evaluate potentially hundreds of viewpoints in each planning iteration. We also compute the mean and variance of the evaluation metrics across 50 steps for 5 different seeds, presented in Fig. 3. We note that our method exhibits significantly lower variance than the baselines. The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
VI Real-world Experiments and Benchmark
Figure 4: A qualitative comparison of the novel views rendered using the map generated by our method, onboard and in real-time, to the baseline and the ground truth. The first row is the rendered color images and the second row is the depth images.
Table II: Quantitative Results of Real-world Experiments.
Methods Env. PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$ Traj. Len. [m] Volume Mappedb [%]
Baseline-Voxel Indoor 1.687 123.9 99.94
Baseline-GS 12.79 0.463 0.481 0.548 0.259 123.9 99.94
[…]
Ours 13.88 0.534 0.442 0.449 0.290 179.1 100.0
Baseline-Voxel Outdoor 0.965 159.9 100.0
Baseline-GS 14.17 0.678 0.406 0.718 0.282 159.9 100.0
[…]
Ours 15.66 0.709 0.368 0.513 0.380 225.2 95.53
a Indoor classes: [floor, chair, table, refrigerator, trashcan]. Outdoor classes: [ground, tree, lamppost, rock, signboard, bench].
b Indoor Volume: 360.0m3. Outdoor Volume: 818.6m3.
VI-A Baselines & Experimental Setup
For a meaningful comparison, we implement a frontier-based exploration algorithm presented in [11] as a baseline method. It uses a 3D voxel map to represent the environment, detect frontiers, and generate plans to the nearest frontier. To improve the performance of the baseline method, we incorporated viewpoint sampling at frontier clusters presented in [27] and used A* to provide a reference path to the frontier for the MPL planner.
We deployed our algorithm and the baseline in 2 different real-world environments for comparison on a Clearpath Jackal robot outfitted with AMD Ryzen 5 3600 and RTX 4000 Ada SFF. In addition, it is equipped with an Ouster OS1 LiDAR for state estimation and Intel Realsense D455 camera for mapping. For the baseline method, we set the voxel resolution to 5cm and limit the frontier cluster size to 10cm. For a fair comparison, the baseline method plans on the voxel map but also maintains a Gaussian map to preserve the quality of view synthesis. To generate the groundtruth data for evaluation, we teleoperated the robot to uniformly sample the environment. The Baseline-Voxel method constructs a voxel map and plans with it. The Baseline-GS method constructs a Gaussian map onboard the robot in addition to the voxel map. We set the budget of Ours-Budget to be the trajectory length of the baseline method. We also carried out experiments for our method with a larger budget.
VI-B Evaluation
Qualitative We evaluate the quality of the Gaussian maps that were constructed in real-time onboard the robot by synthesizing images with the test set poses. Samples of synthesized color images and depth images are presented in Fig. 4. Compared to the Gaussian maps constructed by the baseline method, our synthesized depth images capture the depth value closer to the ground truth images while the color images are sharper with more detail. The rendered views from our method at novel camera poses result in better performance of downstream tasks such as semantic segmentation, details are analyzed in the next section.
Metrics. We computed the same image metrics on the RGB and depth images as in the simulation experiments. We rendered 1000 novel views according to the groundtruth poses from the test set and compared them against the corresponding images. We constrained the budget for each method by the distance traveled. To evaluate the coverage of the environment, we computed the volume mapped from the voxel maps built by each method (in post-process for ours).
Results & Discussion. When the budget is set to the trajectory length of the baseline, our method performs comparably with Baseline-GS. When a larger exploration budget is allowed, the frontier baselines terminate early (since they have explored all available frontiers) but our method is able to significantly improve map reconstruction quality. We attribute this to the information gain metric which enables re-visiting of areas in the environment to further improve map quality. We note that Baseline-Voxel is not able to render RGB images since it does not encode color information and also performs poorly in depth rendering due to the discretization of the voxel map. This further motivates the use of a Gaussian map that contains color in addition to geometric information. Our method optimizes for image rendering quality while still achieving mapped volume comparable with the baseline which prioritizes coverage.
To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also evaluate the rendered images from each method on the task of semantic segmentation. We used Grounded SAM 2 [50, 51, 52, 53, 54] to perform semantic segmentation on the rendered and groundtruth images and computed the mean Intersection over Union (mIoU). Our approach achieves better mIoU scores for both scenes, indicating higher fidelity of the rendered images.
VII Limitations and Future work
In this work, we present a framework for real-time active exploration and mapping compatible with Gaussian Splatting. In future work, we wish to explore the use of Gaussian opacities to consider occluded Gaussians and improve the uncertainty estimates. Another future direction is to consider semantic features together with the Gaussian maps to perform complex tasks in the environment like object search and represent dynamic scenes in our framework. In this work, we use LiDAR to provide odometry but in future work, we aim to formulate this entire framework using a single sensor.

### v2

@@
 Abstract
-We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing information-rich maps.
-Further, we develop a parallelized motion planning algorithm that can exploit the Gaussian map for real-time navigation.
+We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing dense maps.
+Further, we develop a GPU-accelerated motion planning algorithm that can exploit the Gaussian map for real-time navigation.
 The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy.
-We show through simulation experiments that our method is competitive with approaches that use alternate information gain metrics, while being orders of magnitude faster to compute.
-In real-world experiments, our algorithm achieves better map quality (10% higher Peak Signal-to-Noise Ratio (PSNR) and 30% higher geometric reconstruction accuracy) than Gaussian maps constructed by traditional exploration baselines.
+We show through simulation experiments that our method yields comparable Peak Signal-to-Noise Ratio (PSNR) and similar reconstruction error to state-of-the-art approaches, while being orders of magnitude faster to compute.
+In real-world experiments, our algorithm achieves better map quality (at least 0.8dB higher PSNR and more than 16% higher geometric reconstruction accuracy) than maps constructed by a state-of-the-art method, enabling semantic segmentation using off-the-shelf open-set models.
 Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
@@
 1 shows the elements of our approach.
-We use the Gaussian splatting approach proposed in [14] to generate Gaussian maps.
-We propose a novel information gain metric to compute the utility of regions in the environment and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path.
+We use the Gaussian splatting approach proposed in [14] to generate maps.
+We propose an efficient information gain metric and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path.
 Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
@@
 [B] Robot navigating to unobserved areas (right) with high information gain while maximizing information along the trajectory.
-Gaussian frontiers colored cyan (light to dark) in increasing estimated information gain.
+Map regions colored light to dark in increasing estimated information gain.
 In summary, the contributions of this paper are:
   1. 1.
-    A unified framework for online mapping and planning with Gaussians.
+    A unified framework for online mapping and planning built only on a Gaussian splatting-based map, eliminating the need for additional volumetric representations.
   2. 2.
-    An approximate information gain metric that is easy to compute and capable of running in real-time onboard an unmanned ground vehicle.
+    An information gain heuristic that is easy to compute and capable of running in real-time onboard a robot.
   3. 3.
-    A method to compute frontiers in a map built using Gaussian splatting.
-  4. 4.
-    Extensive experiments in both indoor and outdoor environments to evaluate our framework.
+    A real-time exploration system built on Gaussian splatting, comprehensively validated in both simulation and real-world experiments across diverse indoor and outdoor environments.
+We will release the full autonomy framework as open-source.
 II Related Work
-Map Representation.
 To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community.
@@
 In this work, we study the problem of active mapping with the learned map representations.
-Active Simultaneous Localization and Mapping (SLAM).
-The problem of exploration and active SLAM has been widely studied in the past decade.
+The problem of exploration and active Simultaneous Localization and Mapping (SLAM) has been widely studied in the past decade.
 The classical exploration framework uses a model-based approach to actively navigate towards frontiers [11, 12] or waypoints that have the highest information gain [6, 7, 8, 9, 10, 23, 24].
@@
 [36] leverages Fisher information to compute pixel-wise uncertainty on rendered images.
-Where prior work in implicit and radiance field representations necessitates use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a Gaussian map representation encodes physical parameters of the scene, which motivates estimating information gain from the Gaussians directly.
-Navigation in Radiance Fields.
+Where prior work in implicit and radiance field representations use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a map represented by Gaussians encodes physical parameters of the scene, which motivates estimating information from the Gaussian parameters directly.
 Prior work has also considered planning directly in radiance fields.
@@
 [41] use a pre-generated Gaussian map to compute safe polytopes to generate paths and also present a method to synthesize novel views with a coarse localization of the robot and refine the estimate by solving a Perspective-n-Point (PnP) problem.
-The probabilistic representations of free space has also been utilized for motion planning [42] where authors use uncertainties in the learned representation to provide probabilistic guarantees on collision-free paths.
-III Problem Specification
+The probabilistic representations of free space has also been utilized for motion planning [42] where authors use uncertainties in the learned representation to provide guarantees on collision-free paths.
+III Problem Specification & Preliminaries
 Figure 2: The proposed active mapping framework.
 The proposed framework contains two major components, the planning module and the mapping module.
-As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation $m$ at every step and computes the utility of geometrically clustered frontiers.
-The information is then passed to the planning module which comprises the topological graph and motion primitive library ([B]).
-The topological graph adds sampled viewpoints as nodes and passes along a planned path to the trajectory planner.
+As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation at every step and computes the utility of cuboidal regions (Sec.
+IV-B1).
+The utility of each region is then passed to the planning module which comprises the topological graph and motion primitive library ([B]).
+Viewpoints are sampled around regions in the map and added to the topological graph.
+A high-level guidance path is generated from the topological graph to viewpoints with highest utility and passed to the trajectory planner.
 The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper).
 The planned trajectory is executed by the robot to get a new set of observations.
-We have an unknown map of the environment $m^{*}$ that belongs to a space of maps $\mathcal{M}$. Our goal is to solve the following problem
-Problem 1.
-Information-Driven Exploration.
-$\min_{x_{1:T}}$ $\ \mathbb{E}_{x}[\rho(h(x,\hat{m}),h(x,m^{*}))]$ (1)
-$s.t$ $\ \hat{m}=\Phi(y_{1:T},x_{1:T})$
-$\ y_{t}=h(x_{t},m^{*})\ \forall 1\leq t\leq T.$
-In the equation above, the expectation is taken over a random evaluation point $x$.
-The states belong to the state space $\mathcal{X}$, whereas observations belong to the observation space $\mathcal{Y}$.
-The observation function $h:\mathcal{X}\times\mathcal{M}\rightarrow\mathcal{Y}$ takes as input the robot state together with the map of the environment, and outputs the rendered image.
-$T$ represents the exploration budget.
-$m^{*}$ is the true map of the environment.
-$\hat{m}$ is the map of the environment that is estimated from the sequence of measurements $(y_{t})_{1\leq t\leq T}$ taken from vantage points (states) $(x_{t})_{1\leq t\leq T}.$ The discrepancy function $\rho$ captures the difference between the image rendered based on the estimated map, and the actual map.
-The sequence of vantage points is constrained to lie in free space $\mathcal{X}_{free}$.
-Since Prob (1) is parameterized by $m^{*}$, which is a priori unknown, the problem is ill-posed.
-We are instead interested in synthesizing a (near) optimal policy $(\pi_{t})_{t=0}^{T}$, consisting of a sequence of functions that map the history of observations to the optimal action to take at the current step.
-In particular, for any $i\leq T$, the function
-$\pi_{t}:\mathcal{Y}_{1}\times\cdots\times\mathcal{Y}_{t}$ $\rightarrow\mathcal{X}$ (2)
-$\pi_{t}(y_{1},y_{2},...,y_{t})$ $=x_{t}$
-ought to specify the next vantage point given the the history of observations accrued thus far.
-Finding the optimal sequence of policies is a challenging, high-dimensional, optimization problem.
-We therefore develop an approximate method for solving it.
-The approximation lies in a particular, but suitable, choice of representation for $m^{*}$, together with a way of using it for effective exploration.
+Our goal is to construct an estimate $\hat{G}$ of the true map of the environment $G^{*}$, which is a priori unknown.
+The quality of the constructed map $\hat{G}$ is evaluated using a hidden test set $\mathcal{T}$ of tuples of poses and corresponding noise-free measurements, $(x_{test}^{*},y_{test}^{*})_{test\in\mathcal{T}}$.
+We aim to solve the following problem
+$\mathop{\arg\min}\limits_{\hat{G},\ x_{1:T}}$ $\sum_{{test}\in\mathcal{T}}\mathcal{L}(y^{*}_{test},h(x^{*}_{test},\hat{G}))$ (1)
+$s.t$ $\ \hat{G}=\Phi(x_{1:T},y_{1:T}),$
+where $\mathcal{L}(\cdot)$ captures the difference between a synthesized and a true measurement, $\Phi(\cdot)$ is the function that constructs $\hat{G}$ from the set of measurements $y_{1:T}$ obtained at poses $x_{1:T}$, $h(\cdot)$ is the rendering function that synthesizes a measurement given a pose and a map, and $T$ is the time budget for exploration.
+The problem above is ill-posed, as the set of test poses and measurements $\mathcal{T}$ is hidden.
+We approach this challenge by solving the following two problems in a receding horizon scheme.
+At time $k\leq T$, we first solve a mapping problem that involves minimizing the difference between the observed measurements $y_{1:k}$ and synthesized measurements $\hat{y}_{1:k}$:
+$\Phi(x_{1:k},y_{1:k}):=\mathop{\arg\min}\limits_{\hat{G}}\sum_{s=1}^{k}\mathcal{L}(y_{s},h(x_{s},\hat{G})).$ (2)
+Then, given the posterior estimate of the map $\hat{G}$, we find the subsequent viewpoint $x_{k+1}$ that maximizes the mutual information $\mathcal{I}(\cdot)$ between the map $\hat{G}$ and the measurement $y_{k+1}$ observed from the corresponding pose:
+$\max_{x_{k+1}}\ \mathcal{I}(\hat{G};y_{k+1}|x_{k+1}).$ (3)
+Gaussian Splatting.
+We use 3D Gaussian Splatting (3DGS) [5] to represent the environment as a collection of isotropic Gaussians $\hat{G}$.
+Each Gaussian is parameterized with 8 values representing RGB intensities ($c$), 3D position ($\mu$), radius ($r$), and opacity ($\alpha$).
+To generate the measurement at a given camera pose of the estimated map $\hat{G}$, 3D Gaussians are sorted in increasing order of depth relative to the camera pose and are then splatted onto the image plane.
+In particular, each splatted Gaussian will have a projected 2D mean ($\mu_{2D}$) and radius ($r_{2D}$) on the image plane,
+$\mu_{2D}=\frac{KR_{k}^{\intercal}(\mu-T_{k})}{d},\ r_{2D}=\frac{fr}{d},\ d=e_{3}^{\intercal}R_{k}^{\intercal}(\mu-T_{k}),$
+where $K$ is the intrinsic matrix, $R_{k}$ and $T_{k}$ are the rotational and translational components of the camera pose in the world frame at time $k$, $f$ is the focal length, and $e_{3}=[0,0,1]^{\intercal}$.
+For a given pixel $p$ in the image, its color $\mathcal{C(\cdot)}$ and depth $\mathcal{D(\cdot)}$ can be obtained as
+$\mathcal{C}(p)=\sum_{i=1}^{n}c_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p))\$
+and
+$\mathcal{D}(p)=\sum_{i=1}^{n}d_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p)),$
+where
+$f_{i}(p)=\alpha_{i}\exp\left\{-\frac{||p-\mu_{2D,i}||^{2}}{2r_{2D,i}^{2}}\right\}.$
+At every iteration, we render color image $\hat{y}_{c}$ and depth image $\hat{y}_{d}$. We set the function $\mathcal{L}$ (from eq. (2)) to
+$\mathcal{L}$ $=\frac{1}{N_{p}}\sum_{p\in y}\left(|\hat{y}_{d}-y_{d}|+\lambda_{1}|\hat{y}_{c}-y_{c}|\right)$ (4)
+$+\lambda_{2}(1-SSIM(\hat{y}_{c},y_{c}))$
+which is a weighted combination of the L1 loss on depth and rendered pixel colors and the structured similarity index measure (SSIM) to update the parameters of the Gaussians, and $N_{p}$ is the number of pixels in the image.
 IV Method
@@
 The mapping module (Sec.
-IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the utility of Gaussian frontiers.
-The information is then passed to the planning module for planning guidance paths (Sec.
+IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the uncertainty of Gaussians in the map.
+This is then passed to the planning module for planning guidance paths (Sec.
 IV-B1) and trajectories (Sec.
@@
 IV-A Mapping & Uncertainty Estimation
-IV-A1 Mapping
-We adopt the 3D Gaussian Splatting (3DGS) approach to represent the environment as a map of 3D Gaussians.
-3DGS enables real-time radiance field rendering by projecting and blending the 3D Gaussians to render images.
-The parameters of the 3D Gaussians are optimized through mapping iterations to improve the representation of the scene.
 We build the 3DGS mapping module upon the [14] framework.
-The poses of the robot are obtained using [43].
-IV-A2 Gaussian uncertainty for viewpoint selection
-We estimate uncertainty in the map representation directly on the 3D Gaussians.
-We motivate our metric of information gain leveraging the theory of Kalman filtering.
-In particular, the updates of the Kalman filter [44] are given by
-$m_{k|k}$ $=m_{k|k-1}+K(y_{k}-Hx_{k|k-1})$ (3)
-$K$ $=P_{k|k-1}H^{T}S^{-1}$
-$S$ $=HP_{k|k-1}H^{T}+R$
-$P_{k|k}$ $=P_{k|k-1}-KHP_{k|k-1}.$
-where $K$ is the Kalman Gain, $P$ is the covariance of the map, $H$ is the linearized measurement model, $S$ is the innovation covariance and $R$ is the covariance of the measurement noise.
-The following identity holds:
-$Cov_{k|k-1}(m_{k|k}-m_{k|k-1})=P_{k|k-1}-P_{k|k}$ (4)
-The pertinent aspect of the relation above is that the prior covariance of the change in the map parameters upon observing a new measurement is equal to the information gain of the latter.
-Although this argument can be rigorously shown in the context of linear measurement models, we nevertheless use it as a measure of future information gain in the context of measurements with a Gaussian representation of the environment.
-Gaussian uncertainty.
-Following the above motivation, we compute the change in Gaussian parameters across each mapping optimization and use the magnitude of the parameter updates $\|\hat{m}_{t}-\hat{m}_{t-1}\|$ as a measure of uncertainty.
-Gaussians with large changes in parameters represent parts of the scene with high uncertainty and conversely Gaussians with stable parameters have low uncertainty.
-We use this to approximate information gain for viewpoint selection.
-Given a map $m$ represented by a set of Gaussians $G$, we classify the Gaussians in the scene $G=\{G_{H},G_{L},G_{O}\}$ into three disjoint subsets, $G_{H}$ which consists of Gaussians with high uncertainty, $G_{L}$ which consists of Gaussians with low uncertainty and $G_{O}$ for the rest.
-For a Gaussian $g_{i}\in G$ and pose $x_{t}$ at timestep $t$, we define the binary visibility function $v(x_{t},g_{i})$ capturing whether $g_{i}$ is in the field of view of $x_{t}$.
-To maximize information gain while exploring the environment, we aim to find the viewpoint with the highest utility $\xi(x_{t})$ that maximizes the number of high-uncertainty Gaussians and minimizes the number of low-uncertainty Gaussians in the field of view at each timestep.
-The first term in Equ.
-5 encourages observations of parts of the scene with high uncertainty while the second term encourages exploration of unseen areas of the scene.
-The balance of exploitation and exploration is weighted by the term $\lambda_{\xi}$.
-$\xi(x_{t})=\sum_{i\in L}v(x_{t},g_{i})-\lambda_{\xi}\sum_{i\in H}v(x_{t},g_{i})$ (5)
-Gaussian frontiers.
+Similar to [14], we set all Gaussians to be isotropic.
+As in [14], we set $\lambda_{1}=0.4,\lambda_{2}=0.1$.
+At the beginning of each mapping iteration, we initialize new Gaussians at positions corresponding to the current color and depth measurements.
+The parameters of these Gaussians are refined through backpropagation.
+As in [5], we prune Gaussians with low opacity or large radius.
+Uncertainty Estimation. The objective of finding the next-best-view can then be expressed as
+$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathcal{I}(y_{k+1};\hat{G}|x_{k+1}).$ (5)
+The information gain $\mathcal{I}(\cdot)$ is defined as
+$\mathcal{I}$ $(y_{{k+1}};\hat{G}|x_{k+1})$ (6)
+$=\mathbb{H}(y_{k+1}|x_{k+1})-\mathbb{H}(y_{k+1}|\hat{G},x_{k+1})$
+where $\mathbb{H}(\cdot)$ is the entropy.
+We make the following assumptions in this work to approximate the information gain efficiently: (a) $\hat{G}\sim\mathcal{N}(\mu_{\hat{G}},\Sigma_{\hat{G}})$; (b) $y_{k+1}|\hat{G}\sim\mathcal{N}(h(x_{k+1},\hat{G}),\Sigma_{y})$; (c) constant isotropic variance $\Sigma_{y}$.
+From (a) and (b), the information gain can be simplified as
+$\mathcal{I}(y_{{k+1}};\hat{G}|x_{k+1})$ $=\frac{1}{2}\log\det\left(\frac{J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}+\Sigma_{y}}{\Sigma_{y}}\right)$ (7)
+$=\frac{1}{2}\log\det(I+\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal})$
+where $J_{k+1}=\frac{\partial h(x_{k+1},\hat{G})}{\partial G}$. Using the first order approximation of $\log\det$ about the identity, we can define a proxy metric for scoring next best views as
+$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathrm{Tr}(\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}).$ (8)
+As in [36], assumption (c) allows us to simplify the optimization by ignoring the effect of $\Sigma^{-1}_{y}$.
+Evaluating $J_{k+1}$ is computationally expensive because it typically requires rendering at new poses.
+To achieve real-time evaluation of multiple candidate viewpoints, we replace the Jacobian term $J_{k+1}$ with a binary matrix that corresponds to its sparsity pattern.
+Intuitively, this captures the set of Gaussians within the field of view at pose $x_{k+1}$.
+The key insight behind this approximation is that we estimate information gain using the current uncertainty, based on the assumption that local measurements lead to precise local map estimates.
+We show empirically that the exploration performance based on our metric (in Sec.
+V-A) is similar to the performance based on eq.
+(8).
+We propose a heuristic for estimating uncertainty $\Sigma_{\hat{G}}$ based on the magnitude of the change in means ($\mu$) of the Gaussians, associating larger displacements with higher uncertainty.
+The motivation behind this lies in the fact that $\Sigma_{\hat{G}}$ can be approximated with the empirical Fisher Information matrix [43], $\Sigma_{\hat{G}}\approx\nabla_{G}\mathcal{L}(x_{k},y_{k})\nabla_{G}\mathcal{L}(x_{k},y_{k})^{\intercal}$, which in turn can be directly linked to the square of the updates of Gaussian parameters tuned by the gradient descent algorithm.
+In practice, we observe that the square of the magnitude of the updates outlined above is sensitive to measurement noise (in Sec.
+V-A), so we use the magnitude of the change instead.
+Empirically, we observe that the means of Gaussians located near boundaries between observed and unobserved space exhibit significant changes over successive updates, indicating high uncertainty, as visualized in Fig.
+2.
+Consequently, directing exploration toward these high-uncertainty Gaussians is analogous to following frontiers in grid-based exploration, making this heuristic effective for autonomous exploration.
+Importantly, through careful bookkeeping, Gaussians that remain unchanged in the most recent update retain their previously computed displacement values, ensuring that our heuristic provides a consistent measure of each Gaussian’s movement over time.
+In addition to the boundaries between observed and unobserved space, areas within observed space that have insufficient measurements also exhibit relatively high uncertainty.
+This further directs the robot to explore these areas and gather more information, thereby improving the overall map quality.
+IV-B Hierarchical Planning
+In spite of the approximations introduced in the previous section to compute information gain obtained by navigating to a given area in space in a computationally efficient way, the task of finding such a region is still a challenging optimization problem.
+For this reason, we address the latter task through a hierarchical planning framework.
+The high-level planner provides guidance to map particular regions of the environment and a path to that region from the known space in the environment.
+The low-level (trajectory) planner finds a trajectory that is dynamically-feasible (i.e.
+obeys the robot’s physical constraints), collision-free and locally maximizes the information gain along the path.
+IV-B1 High-level planner
 In contrast to traditional mapping representations, the Gaussian map does not encode occluded and free space.
-Instead of frontiers of unobserved regions, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next, since Gaussians that have few observations or lie at the edges of observed regions naturally have higher uncertainty.
-For a region $o\subset m$, we consider the set of Gaussians $\{g\sim\mathcal{N}(\mu,\sigma^{2})\in G\ |\ \mu\in o\}$ with cardinality $N_{g}$ in the region and compute the mean uncertainty of that frontier
-$\Omega=\frac{1}{N_{g}}\sum^{N_{g}}_{i}\|\mu^{t}_{i}-\mu^{t-1}_{i}\|.$ (6)
-This provides a unified framework for estimating information gain to optimize map quality while implicitly identifying frontiers to guide the exploration of unknown regions.
-IV-A3 Traversable region segmentation
-We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor.
-This necessitated the removal of Gaussians on the ground for planning.
-This also fulfilled an important role in reducing the number of Gaussians that have to be queried in the collision check.
-IV-B Hierarchical Planning
-We represent the planning problem as a hierarchical planning problem.
-The first level provides guidance to map particular regions of the environment and a path to that region from the known space in the environment.
-The second level finds a trajectory that is dynamically-feasible (i.e.
-obeys the robot’s physical constraints); collision-free and locally maximizes the information along the path.
-IV-B1 High-level guidance
-We formulate a high-level guidance path that allows us to (i) plan to transition to high uncertainty regions of space which might involve traveling through those with low uncertainty; and (ii) estimate the traversable space in the known space without explicitly having to plan long-range trajectories with computationally expensive collision checks.
-At the higher level of our planner, we construct a graph $\mathcal{G}(N,E)$ by incrementally adding nodes along the traveled path.
-The graph consists of two types of nodes $n_{o},n_{p}$ where $n_{o}$ are observed nodes (along our odometry) and $n_{p}$ are predicted nodes.
+Instead of computing (geometric) frontiers, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next.
+We evenly partition the a priori known enclosing space of the desired map into cuboidal regions.
+At time ${k}$, for a region $o$, denote the cardinality of the Gaussians in the region by $M_{k}$.
+We compute the mean uncertainty of that region $\Omega_{k}=1/{M_{k}}\left(\sum_{\mu\in o}\|\mu_{k}-\mu_{{k-1}}\|_{2}\right).$ We formulate a high-level guidance path that allows us to (i) navigate to regions of high uncertainty; and (ii) utilize the known traversable space to plan long-range trajectories without computationally expensive collision checks.
+At the high level of our planner, we construct a tree by incrementally adding nodes along the traveled path.
+The tree consists of two types of nodes: odometry nodes and viewpoint nodes.
 At each planning iteration, we sample a fixed number of viewpoints around the identified regions.
-We can compute the shortest viewing distance from the Gaussian frontier $f\in\mathbb{R}^{3}$ to the optical center of the camera as
-$d_{1}=\frac{||f^{T}\left[b_{1}\ b_{2}\right]||_{2}}{tan\left(\phi/2\right)}$
-where $b$ is the camera axis and $\phi$ is the field of view of the camera.
-Each sampled viewpoint is then assigned the utility computed by the Gaussian frontiers and connected to the closest point in the graph.
-We then use Dijkstra’s search algorithm to find the shortest path from the current robot location in the graph to all the predicted nodes in the graph.
-We then compute the cost-benefit of a path using $\Omega/e^{d_{2}}$ where $\Omega$ is the estimated information gain and $d_{2}$ is the distance to the node [45].
+We compute the shortest viewing distance from the Gaussian region centroid to the optical center of the camera, given the camera intrinsics.
+Each sampled viewpoint is then assigned the utility computed for that region and connected to the closest odometry node in the tree as a viewpoint node.
+We then use Dijkstra’s algorithm to find the shortest path from the current robot location in the tree to all the viewpoint nodes in the tree.
+Finally, we compute the cost-benefit of a path using $\Omega/e^{d}$ where $\Omega$ is the estimated information gain and $d$ is the distance to the node [44].
 The maximal cost-benefit path is sent to the trajectory planner.
-IV-B2 Trajectory Generation
-The second level planning problem involves finding a collision-free trajectory from start to goal that meets the dynamic constraints of the robot while optimizing the information along the path.
-The problem is defined as follows:
-Problem 2.
-Trajectory Planning. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
-$\min_{u(\cdot),\tau}\ J(u(\cdot),\tau)$ (7)
-$s.t.\ \text{for all }t\in[0,\tau]$
-$\dot{x}(t)=f(x(t),u(t)),\ x(t)\in\mathcal{X}_{free},$
-$||v(t)||_{2}\leq v_{max},\ |\omega(t)|\leq\omega_{max},$
-$x(0),\theta(0)=x_{0},\theta_{0},\ x(\tau)\in\mathcal{X}_{goal}$
-where $J$ is a function that encodes a positive combination of execution time $\tau$ and the control effort $J_{c}(u(\cdot))$.
-We use the unicycle model as the robot dynamics.
+IV-B2 Trajectory planner
+We partition the current map $\hat{G}$ into three disjoint subsets: $\hat{G}_{H}$, which consists of Gaussians with high uncertainty, $\hat{G}_{L}$, which consists of Gaussians with low uncertainty, and $\hat{G}_{O}$ for the rest.
+For a Gaussian $g\in\hat{G}$ and pose $x$, we define the binary visibility function $v(x,g)$ capturing whether $g$ is in the field of view of $x$.
+To avoid rendering each view to check for occlusions, we cull the Gaussians that are beyond the perception range of the sensor.
+To maximize information gain while exploring the environment, we aim to obtain viewpoints with high utility $\xi$ that maximize the number of high-uncertainty Gaussians and minimize the number of low-uncertainty Gaussians in the field of view.
+Trajectories are evaluated based on the sum of the utilities of the viewpoints in each trajectory given by
+$\xi(x,\hat{G})=\sum_{g\in\hat{G}_{H}}v(x,g)-\lambda_{\xi}\sum_{g\in\hat{G}_{L}}v(x,g).$ (9)
+The first term in eq.
+(9) encourages additional observations of parts of the scene with high uncertainty.
+The second term penalizes observation of stable Gaussians in explored areas and encourages exploration of unseen parts of the environment.
+The two terms are weighted by the $\lambda_{\xi}$.
+Given the path generated by the high-level planner, we select the furthest point $p_{goal}$ along the path that lies within a local planning range $R_{H}$, and set it as the center of the goal region $\mathcal{X}_{goal}$.
+We use the unicycle model as the robot dynamics $f(\cdot)$.
 The robot state $x=[p,\theta]\in\mathbb{R}^{2}\times S^{1}$ consists of its position ($p$) and heading ($\theta$).
 The control inputs $u=[v,\omega]\in\mathbb{R}^{2}$ consist of linear velocity ($v$) and angular velocity $(\omega)$.
-Inspired by [46], we solve problem 2 by performing a tree search on the motion primitives tree.
-Motion primitives are generated with fixed control inputs over a time interval $\delta\tau$ with known initial states.
+We solve the trajectory planning problem in 2 steps: (i) finding feasible, collision-free trajectory candidates, and (ii) selecting a trajectory that maximizes the information gain.
+The first problem is defined as follows:
+Problem 1.
+Trajectory Candidate Generation. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
+[…]
+$s.t.\ \forall t\in[0,\tau],\ x(0)=x_{0},\ x(\tau)\in\mathcal{X}_{goal},$
+$\dot{x}(t)=f(x(t),u(t)),\ x(t)\in\mathcal{X}_{free},$
+$||v(t)||_{2}\leq v_{max},\ |\omega(t)|\leq\omega_{max}$
+where $\lambda_{t}$ weights the time cost with the control effort, $v_{max}$ and $\omega_{max}$ are actuation constraints, and $x_{0}$ is the initial state.
+Motion Primitive Tree Generation.
+Inspired by [45], we solve problem 1 by performing a tree search on the motion primitives tree.
+Motion primitives are generated with fixed control inputs over a time interval with known initial states.
 Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $[0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs.
-Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization $d\tau$.
+Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization.
+Collision Checking.
 We sample a fixed number of points on each motion primitive to conduct collision checks.
-Since we have an uncertain map, we relax $x\in X_{free}$, to a chance constraint.
-First, we measure the distance from the discretized positions
-$w\sim\mathcal{N}(\kappa,r_{robot}^{2})\in W|\kappa=x_{t}$
-along the search path to all Gaussians
-$\{g\sim\mathcal{N}(\mu,\sigma^{2})\in G\}$
-in the map. Our constraint then boils down to the probability of the distance between the test point and the set of Gaussians in the scene
-$d_{g}=\inf||W-G||$
-being less than 0 i.e.
-$P(d_{g}<0)\leq\eta$ where $\eta$ is the tolerable risk.
-When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $\sigma$ of each Gaussian is scaled by a factor of $\lambda_{g}$.
-The truth value of the test is determined by comparing the distance to the closest Gaussian with $r_{robot}+\lambda_{g}\sigma$, as illustrated in Fig.
+Since we have an uncertain map, we relax $x\in X_{free}$ to a chance constraint.
+Let $d_{\kappa}$ be the minimum distance between a test point and the set of Gaussians.
+Our constraint then amounts to the probability of the distance between the test point and the set of Gaussians in the scene being less than some allowed tolerance $\gamma$ with probability $\eta$ i.e.
+$P(d_{\kappa}<\gamma)\leq\eta$.
+When checking for collisions, each sampled point is bounded with a sphere of radius $r_{robot}$ and the radius $r$ of each Gaussian is scaled by a factor of $\lambda_{g}$.
+The truth value of the test is determined by comparing the distance to all Gaussians with $\gamma:=r_{robot}+\lambda_{g}r$, as illustrated in Fig.
 2.
-The rationale for the latter test is as follows.
-Assuming that the closest obstacle is unique, the probability of collision with some obstacle is well-approximated by the probability of collision with the nearest obstacle.
-This holds by virtue of the exponential decay of the Gaussian distribution with the distance from the origin.
-As a result, specifying an upper bound on the probability of collision is approximately the same as specifying the probability of collision with the closest obstacle.
-We develop a GPU-accelerated approach for testing all sampled points with all Gaussians from the map at once while growing the search tree.
+Note that setting $\lambda_{g}=3$ is equivalent to the test proposed by [37].
+However, unlike the continuous trajectory optimization approach in [37], which requires subsampling to provide soft constraints, our search-based method allows dense checks.
+We are implicitly assuming that 3DGS accurately captures geometry of the scene and we notice this holds empirically since we initialize Gaussians in the map using the depth measurements from the RGB-D sensor.
+In the case of collision checking against non-isotropic Gaussians, the principal axis of the Gaussian can be taken as the radius.
+We develop a GPU-accelerated approach for testing collisions of all sampled points with all Gaussians from the map at once while growing the search tree.
 This allows the real-time expansion of the search tree.
-The cost of each valid motion primitive is defined as
-$J(u(\cdot),\tau):=\lambda_{t}\tau+(v^{2}+w^{2})\tau,$
-where $\lambda_{t}$ weights the time cost with the control efforts. Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as
-$h(p):=||p_{goal}-p||_{2}/v_{max}.$
-We use A* to search through the motion primitives tree and keep top $k$ candidate trajectories for the subsequent information maximization.
-The information of each trajectory is evaluated with state sequence $[x(\delta\tau),\ldots,x(\tau)]$ on the trajectory according to (5).
-Finally, the trajectory with the highest information $\left(\sum_{t=0}^{\tau}\xi(x(t))\right)$ is selected and executed by the robot.
+Tree Search.
+The cost of each valid motion primitive is defined by Prob.
+1.
+Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as $h(p):=||p_{goal}-p||_{2}/v_{max}$.
+We use A* to search through the motion primitives tree and keep top $N_{traj}$ candidate trajectories for the information gain maximization.
+Information Gain Maximization.
+A state sequence $\{x_{i}({l\tau/L})\ |\ l\in\{0,\ldots,L\}\}$ containing the end state of each of $L$ segments in the $i$-th trajectory is used for evaluating the information along the trajectory.
+The information of each state is evaluated according to eq.
+(9).
+Finally, the trajectory with the highest information
+$\mathop{\arg\max}\limits_{i\in\{1,\ldots,N_{traj}\}}\sum_{l=0}^{L}\xi\left(x_{i}\left(\frac{l\tau}{L}\right)\right),\ l\in\{0,\ldots,L\}$ (11)
+is selected and executed by the robot.
 IV-C Implementation Details
-We implemented the proposed method using PyTorch2.4.
-Most of our parameters of the Gaussian map were set to the default configuration of [14] for the TUM dataset.
+We implemented the proposed method using PyTorch2.4 and the mapping module following [14].
+We note that our framework is compatible with other Gaussian splatting approaches if they can meet the compute and latency requirements.
+The parameters of the mapper were set to the default configuration of [14] for the TUM dataset.
 For simulation experiments, we used 50 mapping iterations.
 For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence.
-The mapping module ran at 1Hz and the planning module ran at 3Hz onboard the robot.
-We set Gaussian frontier size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$.
-V Simulation Results & Ablations
-Table I: Simulation Results on 120 iTHOR scenes with 20 steps
-Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$ t/step $\downarrow$
-[dB] [m] [s]
-Ensemble 12.589 0.483 0.612 1.064 11.094
-FisherRF 13.702 0.558 0.549 1.016 5.347
-Ours 13.591 0.533 0.576 0.975 0.046
-Environments.
-We conduct simulation experiments with the AI2-THOR simulator [47] on the iTHOR dataset.
+The mapping and planning modules are both set to 1Hz onboard the robot.
+For trajectory viewpoint evaluation, we set the threshold for high and low uncertainty Gaussians at half a standard deviation above and below the average magnitude of parameter updates, respectively.
+We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor.
+This necessitated the removal of Gaussians on the ground plane for planning.
+We set the Gaussian region size to 2.5m, $\lambda_{\xi}=1$ and $\lambda_{g}=3$.
+For planning, we set $N_{v}=3$, $v_{max}=0.6$, $N_{\omega}=5$, $\omega_{max}=0.9$, and time discretization 1s.
+V Results
+V-A Simulation Experiments & Ablations
+Table I: Benchmark experiments on iTHOR test scenes with 100 steps
+Methods PSNR SSIM LPIPS RMSE t/step
+[dB] $\uparrow$ $\uparrow$ $\downarrow$ [m] $\downarrow$ [s] $\downarrow$
+Ensemble 21.489 0.762 0.360 0.350 2.667
+FisherRF 24.171 0.849 0.270 0.285 1.299
+RTGuIDE 22.946 0.820 0.305 0.343 0.013
+Table II: Ablation experiments on iTHOR test scenes with 100 steps
+Methods PSNR SSIM LPIPS RMSE t/step
+[dB] $\uparrow$ $\uparrow$ $\downarrow$ [m] $\downarrow$ [s] $\downarrow$
+RTGuIDE 22.946 0.820 0.305 0.343 0.013
+RTGuIDE-sum 20.279 0.740 0.371 0.512 0.007
+RTGuIDE-sq 22.489 0.802 0.308 0.343 0.014
+RTGuIDE w. noise 22.541 0.801 0.329 0.344 0.013
+RTGuIDE-sq w. noise 21.445 0.779 0.351 0.386 0.013
+V-A1 Experiment setup & Baselines
+To evaluate our proposed heuristic, we conduct simulation experiments with the AI2-THOR simulator [46] on the iTHOR dataset.
 The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms.
-Evaluation metrics & Baselines.
-For each scene, we generate a test set of images and evaluate the rendered images from each method against the test set.
+Each step in the simulator involves a translation and rotation.
+The agent is allowed to translate 0.5m in both $x$ and $y$ positions.
+At each allowed position, we sample a set of candidate yaw angles to form a discrete set of possible orientations.
+These candidate positions, along with their corresponding yaw angles, define the full set of potential viewpoints.
+Each viewpoint is evaluated using an uncertainty metric, and the one with the highest uncertainty is selected as the next action.
+We ran experiments on all test scenes from iTHOR with 100 steps each.
+The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
+For each scene, we synthesized novel views from the generated maps from a set of uniformly sampled test poses and computed the image metrics against the ground truth images.
 We evaluate the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [47], Learned Perceptual Image Patch Similarity (LPIPS) [48] on the RGB images and Root Mean-Square-Error (RMSE) on the depth images.
-We evaluate the proposed information gain metric against two other view selection methods.
+We evaluate the proposed method (denoted as Ours) on viewpoints following eq.
+(9) against two other view selection methods.
+For FisherRF, we follow the original implementation to evaluate viewpoints based on the Fisher information.
 For the Ensemble baseline, we train an ensemble of $5$ models with the leave-one-out procedure at each step.
 We compute the patch-wise variance across the rendered images for each sampled viewpoint and ensemble model, and select the viewpoint with the highest variance as the next action.
-For FisherRF, we follow the original implementation to evaluate viewpoints based on the Fisher information.
-Experiments and results.
-In our simulation experiments, we set the grid size of reachable positions to 0.25m.
-At each step, we sample viewpoints with different yaw angles around the current position of the agent as possible actions to take.
-These viewpoints are then evaluated with the corresponding uncertainty metric and the viewpoint with the highest uncertainty is selected as the next action.
-We ran experiments on all 120 scenes from iTHOR with 20 steps each.
-We synthesized novel views according to the test set poses and computed the image metrics against the test set images.
-The averaged results are presented in Tab. I. Our information gain metric performs comparably with FisherRF and
-Figure 3: Comparison of viewpoint selection methods on environment FloorPlan7 of the iTHOR dataset with 50 steps.
-The average and variance are computed across 5 different seeds.
-Our proposed method achieves comparable performance to the other methods while exhibiting significantly smaller variance across runs.
-outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time.
+V-A2 Benchmark experiments
+The averaged results over all test scenes are presented in Tab.
+I.
+Our heuristic performs comparably with FisherRF and outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time.
 Since we are computing the uncertainty directly on the Gaussians instead of the image, we avoid the computation cost of rendering every sampled viewpoint.
 This efficiency in computation is crucial in a sampling-based planner where we evaluate potentially hundreds of viewpoints in each planning iteration.
-We also compute the mean and variance of the evaluation metrics across 50 steps for 5 different seeds, presented in Fig.
-3.
-We note that our method exhibits significantly lower variance than the baselines.
-The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
-VI Real-world Experiments and Benchmark
-Figure 4: A qualitative comparison of the novel views rendered using the map generated by our method, onboard and in real-time, to the baseline and the ground truth.
-The first row is the rendered color images and the second row is the depth images.
-Table II: Quantitative Results of Real-world Experiments.
-Methods Env. PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$ Traj. Len. [m] Volume Mappedb [%]
-Baseline-Voxel Indoor 1.687 123.9 99.94
-Baseline-GS 12.79 0.463 0.481 0.548 0.259 123.9 99.94
+V-A3 Ablation experiments
+For ablation experiments, we implemented RTGuIDE-sum, which solely considers the sum of the uncertainty of Gaussians in the camera view.
+RTGuIDE-sq evaluates the uncertainty of Gaussians as the squared L2 norm following the original derivation in eq.
+(8) and using eq.
+(9) to evaluate the uncertainty of viewpoints.
+To simulate real-world noisy measurements, we added Gaussian noise $(\mu=0,\sigma^{2}=0.1)$ to depth measurements in the simulation to evaluate RTGuIDE w.
+noise and RTGuIDE-sq w.
+noise.
+As shown in Tab.
+II, the ablation experiments on different metrics demonstrate that our proposed approach outperforms the simple uncertainty summation (RTGuIDE-sum), which focuses purely on exploitation.
+Furthermore, the results show that both the squared (RTGuIDE-sq) and standard L2 norms (RTGuIDE) achieve comparable performance when perfect depth measurements are available.
+However, when noise is introduced, the experiments confirm that our chosen metric is more robust and better suited for real-world scenarios where measurement noise is inevitable.
+V-B Planner Performance
+We evaluate the necessity of our proposed GPU-based planning approach in enabling real-time planning.
+In particular, we conducted experiments on trajectory planning with a 5-meter horizon, evaluating both GPU-accelerated collision checking and a fully CPU-based implementation.
+As illustrated in Fig.
+3, the planning time of CPU-based planner increases significantly with the number of Gaussians.
+For example, a 40m $\times$ 40m outdoor parking lot can contain around $3\times 10^{6}$ Gaussians.
+In this case, the CPU-based planner requires 11.65 seconds, whereas the GPU-based planner takes 0.62 seconds, achieving 18$\times$ speedup.
+The GPU-based collision testing enables parallel growth of the search tree at each layer, along with simultaneous and efficient collision checks between all test points and Gaussians directly on the GPU.
 […]
-Ours 13.88 0.534 0.442 0.449 0.290 179.1 100.0
-Baseline-Voxel Outdoor 0.965 159.9 100.0
-Baseline-GS 14.17 0.678 0.406 0.718 0.282 159.9 100.0
+Figure 3: Time spent to plan a trajectory with a 5m horizon (including 129 collision checking points) versus the number of Gaussians in the map.
+V-C Real-world Experiments & Benchmarks
+Table III: Quantitative Results of Real-world Experiments.
+Methods Env. Budget (Time) [s] PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$
+Ensemble Indoor 1 72m2 300 6.27 0.204 0.839 2.051 0.071
+FisherRF 16.80 0.665 0.393 0.242 0.314
+RTGuIDE 17.83 0.737 0.334 0.202 0.338
 […]
-Ours 15.66 0.709 0.368 0.513 0.380 225.2 95.53
-a Indoor classes: [floor, chair, table, refrigerator, trashcan]. Outdoor classes: [ground, tree, lamppost, rock, signboard, bench].
-b Indoor Volume: 360.0m3. Outdoor Volume: 818.6m3.
-VI-A Baselines & Experimental Setup
-For a meaningful comparison, we implement a frontier-based exploration algorithm presented in [11] as a baseline method.
-It uses a 3D voxel map to represent the environment, detect frontiers, and generate plans to the nearest frontier.
-To improve the performance of the baseline method, we incorporated viewpoint sampling at frontier clusters presented in [27] and used A* to provide a reference path to the frontier for the MPL planner.
-We deployed our algorithm and the baseline in 2 different real-world environments for comparison on a Clearpath Jackal robot outfitted with AMD Ryzen 5 3600 and RTX 4000 Ada SFF.
-In addition, it is equipped with an Ouster OS1 LiDAR for state estimation and Intel Realsense D455 camera for mapping.
-For the baseline method, we set the voxel resolution to 5cm and limit the frontier cluster size to 10cm.
-For a fair comparison, the baseline method plans on the voxel map but also maintains a Gaussian map to preserve the quality of view synthesis.
+Ensemble Indoor 2 189m2 600 10.41 0.269 0.787 2.410 0.156
+FisherRF 14.88 0.588 0.451 1.170 0.248
+RTGuIDE 16.40 0.711 0.299 0.689 0.391
+[…]
+Ensemble Outdoor 1 252m2 600 10.59 0.392 0.637 1.705 0.549
+FisherRF 17.58 0.687 0.349 0.519 0.600
+RTGuIDE 20.00 0.744 0.296 0.369 0.622
+[…]
+Ensemble Outdoor 2 480m2 900 16.79 0.538 0.576 1.422 0.159
+FisherRF 19.44 0.680 0.439 0.773 0.177
+RTGuIDE 20.27 0.750 0.345 0.399 0.174
+[…]
+a Indoor 1 classes: [floor, chair, table, refrigerator, cabinet, backpack, plants].
+Indoor 2 classes: [floor, chair, table, couch, cushion, trashcan, television, plants].
+Outdoor 1 classes: [pavement, plants, barrel, traffic cone].
+Outdoor 2 classes: [pavement, grass, tree, bench, lamppost].
+V-C1 Experiment setup & Baselines
+To evaluate the effectiveness of our proposed framework, we compared it with two baselines onboard the robot.
+We first implemented FisherRF [36] by constructing both a voxel map and a Gaussian splatting map online.
+We detect and cluster frontiers from the voxel map and use A* to generate reference paths to frontier clusters.
+Camera views along each path are evaluated to compute information gain.
+To generate collision-free and dynamically feasible trajectories, we generate trajectories along waypoints in the path through MPL [45] in the voxel map considering the unicycle model for robot dynamics.
+We also implemented an Ensemble method and deployed it on the robot, similar to the framework of FisherRF.
+In the ensemble approach, uncertainty is computed as described in Sec.
+V-A1.
+We deployed our method and the baselines in four different real-world environments on a Clearpath Jackal robot outfitted with an AMD Ryzen 5 3600 and RTX 4000 Ada SFF.
+In addition, the platform is equipped with an Ouster OS1 LiDAR for state estimation and a ZED 2i stereo camera for mapping.
+We used [49] to provide odometry and truncated the depth measurements at 5 meters for all methods in all of our experiments.
+For the baseline methods, we set the voxel resolution to 5cm.
 To generate the ground truth data for evaluation, we teleoperated the robot to uniformly sample the environment.
-The Baseline-Voxel method constructs a voxel map and plans with it.
-The Baseline-GS method constructs a Gaussian map onboard the robot in addition to the voxel map.
-We set the budget of Ours-Budget to be the trajectory length of the baseline method.
-We also carried out experiments for our method with a larger budget.
-VI-B Evaluation
-Qualitative We evaluate the quality of the Gaussian maps that were constructed in real-time onboard the robot by synthesizing images with the test set poses.
-Samples of synthesized color images and depth images are presented in Fig.
-4.
-Compared to the Gaussian maps constructed by the baseline method, our synthesized depth images capture the depth value closer to the ground truth images while the color images are sharper with more detail.
-The rendered views from our method at novel camera poses result in better performance of downstream tasks such as semantic segmentation, details are analyzed in the next section.
-Metrics.
-We computed the same image metrics on the RGB and depth images as in the simulation experiments.
-We rendered 1000 novel views according to the groundtruth poses from the test set and compared them against the corresponding images.
-We constrained the budget for each method by the distance traveled.
-To evaluate the coverage of the environment, we computed the volume mapped from the voxel maps built by each method (in post-process for ours).
-Results & Discussion.
-When the budget is set to the trajectory length of the baseline, our method performs comparably with Baseline-GS.
-When a larger exploration budget is allowed, the frontier baselines terminate early (since they have explored all available frontiers) but our method is able to significantly improve map reconstruction quality.
-We attribute this to the information gain metric which enables re-visiting of areas in the environment to further improve map quality.
-We note that Baseline-Voxel is not able to render RGB images since it does not encode color information and also performs poorly in depth rendering due to the discretization of the voxel map.
-This further motivates the use of a Gaussian map that contains color in addition to geometric information.
-Our method optimizes for image rendering quality while still achieving mapped volume comparable with the baseline which prioritizes coverage.
-To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also evaluate the rendered images from each method on the task of semantic segmentation.
-We used Grounded SAM 2 [50, 51, 52, 53, 54] to perform semantic segmentation on the rendered and groundtruth images and computed the mean Intersection over Union (mIoU).
-Our approach achieves better mIoU scores for both scenes, indicating higher fidelity of the rendered images.
-VII Limitations and Future work
-In this work, we present a framework for real-time active exploration and mapping compatible with Gaussian Splatting.
+The exploration budget for all methods is set to $\sim$5x the time needed to fully survey the environment with teleoperation.
+We rendered 500 novel views according to the ground truth poses and compared them against the corresponding images.
+We computed the same image metrics on the rendered RGB and depth images as in the simulation experiments.
+To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also performed an evaluation on the task of semantic segmentation.
+We used Grounded SAM 2 [50] to obtain segmentation masks of the rendered and groundtruth images and computed the mean Intersection over Union (mIoU).
+V-C2 Qualitative results
+Figure 4: Visualization of onboard constructed map
+Figure 5: Qualitative results
+We evaluated the quality of the Gaussian splatting maps that were constructed in real-time onboard the robot by rendering novel views with a set of test poses.
+A visualization of the onboard constructed map is shown in Fig.
+4, and examples of rendered color images are presented in Fig.
+5.
+Constructed in real time onboard the robot, our Gaussian splatting map provides a detailed representation of the environment.
+Being able to render photorealistic images from the map benefits downstream tasks such as semantic segmentation, which are analyzed in detail in the next section.
+V-C3 Quantitative results
+As shown in Tab.
+III, in real-world scenarios when the operational budget is limited, our proposed framework constructs maps of higher quality than the baselines onboard the robot.
+To validate that our approach generates maps of reasonable quality, we also present maps generated with the ground truth samples.
+Results on novel view image rendering show that the maps constructed with our approach achieve 0.8-2.4 higher PSNR than FisherRF and 3.5-11.6 higher PSNR than Ensemble.
+Furthermore, when evaluated on depth image recovery, our method reduces the RMSE of the rendered depth by 16.5%-48% compared to FisherRF, and by 71.4%-90.2% compared to Ensemble.
+These results reflect that our method outperforms the baselines in achieving good coverage of the environments while also optimizing for map quality during exploration.
+For semantic segmentation, our approach achieves better mIoU scores for most experiments, indicating higher fidelity of the rendered images compared to the baselines.
+V-C4 Discussion
+We attribute the performance of our approach to two key factors.
+First, the use of an information metric that is efficient to compute onboard the robot enables smooth and continuous operation.
+Second, since our method does not rely on frontiers extracted from the voxel map for geometric coverage, it allows for revisiting of areas in the environment to further improve map quality.
+VI Limitations and Future work
+In this work, we present a framework for real-time active exploration and mapping with Gaussian Splatting.
 In future work, we wish to explore the use of Gaussian opacities to consider occluded Gaussians and improve the uncertainty estimates.
 Another future direction is to consider semantic features together with the Gaussian maps to perform complex tasks in the environment like object search and represent dynamic scenes in our framework.
-In this work, we use LiDAR to provide odometry but in future work, we aim to formulate this entire framework using a single sensor.
+In this work, we study the active mapping problem which assumes perfect state estimation.
+In future work, we aim to incorporate the state estimation uncertainty and formulate this entire framework using a single sensor.

### v3

@@
 The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy.
-We show through simulation experiments that our method yields comparable Peak Signal-to-Noise Ratio (PSNR) and similar reconstruction error to state-of-the-art approaches, while being orders of magnitude faster to compute.
-In real-world experiments, our algorithm achieves better map quality (at least 0.8dB higher PSNR and more than 16% higher geometric reconstruction accuracy) than maps constructed by a state-of-the-art method, enabling semantic segmentation using off-the-shelf open-set models.
+We show through viewpoint selection experiments that our method yields comparable Peak Signal-to-Noise Ratio (PSNR) and similar reconstruction error to state-of-the-art approaches, while being orders of magnitude faster to compute.
+In closed-loop physics-based simulation and real-world experiments, our algorithm achieves better map quality (at least 0.8dB higher PSNR and more than 16% higher geometric reconstruction accuracy) than maps constructed by a state-of-the-art method, enabling semantic segmentation using off-the-shelf open-set models.
 Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
@@
 Active mapping is a problem of optimizing the trajectory of an autonomous robot in an unknown environment to construct an informative map in real-time.
-It is a critical component of numerous real-world applications such as precision agriculture [1], infrastructure inspection [2], and search and rescue [3] missions.
+It is a critical component of numerous real-world applications such as precision agriculture, infrastructure inspection, and search and rescue missions.
 While nearly all tasks rely on recovering accurate metric information to enable path planning, many also require more fine-grained information.
 Recent advances in learned map representations from the computer vision and graphics communities [1, 2] have opened up new possibilities for active mapping and exploration while maintaining both geometrically and visually accurate digital twins of the environment.
-While prior work effectively solves the problem of information-driven exploration [6, 7, 8, 9, 10] or frontier-based exploration [11, 12, 13], in this work we consider the additional problem of generating radiance fields while also performing autonomous navigation.
-Prior work has also proposed information metrics using novel learned scene representations that are capable of high quality visual reconstruction but are incapable of running in real-time onboard a robot.
+While prior work effectively solves the problem of information-driven exploration [3, 4] or frontier-based exploration [5, 6], in this work we consider the additional problem of generating radiance fields while also performing autonomous navigation.
+Prior work has also proposed information metrics using novel learned scene representations that are capable of high quality visual reconstruction but are incapable of running in real-time onboard a robot [7].
 To enable efficient mapping and planning in these novel representations, we consider approximation techniques for computing the information gain.
@@
 1 shows the elements of our approach.
-We use the Gaussian splatting approach proposed in [14] to generate maps.
+We use the Gaussian splatting approach proposed in SplaTAM [8] to generate maps.
 We propose an efficient information gain metric and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path.
+We utilize 3DGS as a unified representation for both mapping and planning.
+We avoid the cost of maintaining additional volumetric representations – achieving the same granularity of a 3DGS map with a voxel map would require around 19.2 GB of RAM for a $40m\times 40m\times 3m$ space at 1 $cm$ resolution.
+With our GPU-based trajectory planner, we achieve an 18x speedup compared to a CPU-based implementation, as detailed in Sec.
+V-B.
+This allows us to generate collision-free trajectories on millions of Gaussians efficiently.
 Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
@@
   1. 1.
-    A unified framework for online mapping and planning built only on a Gaussian splatting-based map, eliminating the need for additional volumetric representations.
+    A unified framework for online mapping and planning built only on Gaussian splatting, eliminating the need for maintaining additional volumetric representations.
   2. 2.
@@
 II Related Work
+Map Representation.
 To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community.
 The most intuitive but effective volumetric representation has been widely used.
-Voxel-based representation could maintain information such as occupancy [15] or signed distance [16, 17].
-With the recent application of semantic segmentation, semantic maps that contain actionable information have been proposed [18, 19, 20, 21, 22].
-With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [4] and 3D Gaussian Splatting (3DGS) [5] have become popular representations for robotic motion planning.
-In this work, we study the problem of active mapping with the learned map representations.
+Voxel-based representation could maintain information such as occupancy [9] or signed distance [10].
+With the recent application of semantic segmentation, semantic maps that contain additional task-relevant information have been proposed [11, 12, 13].
+With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [1] and 3D Gaussian Splatting (3DGS) [2] have become popular representations for robotic motion planning [7, 14, 15].
+In this work, we study the problem of active mapping with the 3DGS representation.
+Active Mapping.
 The problem of exploration and active Simultaneous Localization and Mapping (SLAM) has been widely studied in the past decade.
-The classical exploration framework uses a model-based approach to actively navigate towards frontiers [11, 12] or waypoints that have the highest information gain [6, 7, 8, 9, 10, 23, 24].
-Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [25, 26, 27, 28].
+The classical exploration framework uses a model-based approach to actively navigate towards frontiers [5] or waypoints that have the highest information gain [3, 4, 7].
+Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [16, 17, 18].
 However, most of the existing work developed their approaches based on classical map representations.
-For example, frontiers are typically defined in voxel maps, information metrics are typically associated with occupancy maps or Signed Distance Field (SDF) maps.
-In this work, we instead consider the active mapping problem with a learned map representation.
-Bayesian neural networks [29, 30] and deep ensembles [24, 31] are common approaches for estimating uncertainty in learned representations.
-[32, 33] use the idea of computing the uncertainty based on consecutive measurements from a moving robot in contrast to train multiple models in the traditional ensemble frameworks.
+In this work, we instead consider the active mapping problem with a 3DGS representation.
+Bayesian neural networks [19, 20] and deep ensembles [7] are common approaches for estimating uncertainty in learned representations.
 Radiance field representations provide additional possibilities for estimating uncertainty through the volumetric rendering process [21].
-[35] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty.
-[36] leverages Fisher information to compute pixel-wise uncertainty on rendered images.
-Where prior work in implicit and radiance field representations use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a map represented by Gaussians encodes physical parameters of the scene, which motivates estimating information from the Gaussian parameters directly.
+NARUTO [22] uses an uncertainty-learning module to quantify uncertainty for active reconstruction.
+CG-SLAM [23] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty.
+FisherRF [24] leverages Fisher information to compute pixel-wise uncertainty on rendered images.
+Where prior work in implicit and radiance field representations use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a map represented by Gaussians encodes physical parameters of the scene, which motivates estimating information from the Gaussian parameters directly without rendering.
+Planning in Radiance Fields.
 Prior work has also considered planning directly in radiance fields.
-[37] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration.
-Sim-to-real approaches leverage the rendering quality of the learnt radiance fields to train downstream tasks such as visual localization, imitation learning [38] by effectively utilizing the representation as a simulator.
-[39] use a liquid neural network to train imitation policies for a quadrotor robot in a Gaussian splat.
+GS-planner [15] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration.
+Sim-to-real approaches leverage radiance fields as a simulator for imitation learning [25].
+Adamkiewicz et al.
 [26] utilize the geometric fidelity of the representation to first map the environment and then perform trajectory optimization to plan paths in these environments.
-[41] use a pre-generated Gaussian map to compute safe polytopes to generate paths and also present a method to synthesize novel views with a coarse localization of the robot and refine the estimate by solving a Perspective-n-Point (PnP) problem.
+Splat-nav [14] uses a pre-generated Gaussian map to compute safe polytopes to generate paths.
 The probabilistic representations of free space has also been utilized for motion planning [27] where authors use uncertainties in the learned representation to provide guarantees on collision-free paths.
+While prior work considers planning in radiance fields, they typically maintain a separate volumetric representation for planning which requires additional memory usage onboard the robot.
+On the other hand, maintaining a secondary representation may lead to inconsistencies in collision checks for planning.
+Existing work that performs collision checking with 3DGS requires subsampling to provide soft constraints.
+In this work, our proposed planning module performs dense checks directly against all Gaussians for collision avoidance.
 III Problem Specification & Preliminaries
@@
 The utility of each region is then passed to the planning module which comprises the topological graph and motion primitive library ([B]).
-Viewpoints are sampled around regions in the map and added to the topological graph.
-A high-level guidance path is generated from the topological graph to viewpoints with highest utility and passed to the trajectory planner.
 The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper).
@@
 We aim to solve the following problem
-$\mathop{\arg\min}\limits_{\hat{G},\ x_{1:T}}$ $\sum_{{test}\in\mathcal{T}}\mathcal{L}(y^{*}_{test},h(x^{*}_{test},\hat{G}))$ (1)
+$\mathop{\arg\min}\limits_{\hat{G},\ x_{1:T}}$ $)\displaystyle\sum_{{test}\in\mathcal{T}}\mathcal{L}(y^{*}_{test},h(x^{*}_{test},\hat{G}$)) (1)
 $s.t$ $\ \hat{G}=\Phi(x_{1:T},y_{1:T}),$
@@
 At time $k\leq T$, we first solve a mapping problem that involves minimizing the difference between the observed measurements $y_{1:k}$ and synthesized measurements $\hat{y}_{1:k}$:
-$\Phi(x_{1:k},y_{1:k}):=\mathop{\arg\min}\limits_{\hat{G}}\sum_{s=1}^{k}\mathcal{L}(y_{s},h(x_{s},\hat{G})).$ (2)
-Then, given the posterior estimate of the map $\hat{G}$, we find the subsequent viewpoint $x_{k+1}$ that maximizes the mutual information $\mathcal{I}(\cdot)$ between the map $\hat{G}$ and the measurement $y_{k+1}$ observed from the corresponding pose:
-$\max_{x_{k+1}}\ \mathcal{I}(\hat{G};y_{k+1}|x_{k+1}).$ (3)
+$).\Phi(x_{1:k},y_{1:k}):=\mathop{\arg\min}\limits_{\hat{G}}\sum_{s=1}^{k}\mathcal{L}(y_{s},h(x_{s},\hat{G}$)). (2)
+Given the posterior estimate of the map $\hat{G}$, we find the subsequent viewpoint $x_{k+1}$ that maximizes the information gain $\mathcal{I}(\cdot)$ given by the mutual information between the map $\hat{G}$ and the measurement $y_{k+1}$ at the corresponding pose:
+$\max_{x_{k+1}}\ \mathcal{I}(\hat{G};y_{k+1}\mid x_{k+1}).$ (3)
 Gaussian Splatting.
@@
 In particular, each splatted Gaussian will have a projected 2D mean ($\mu_{2D}$) and radius ($r_{2D}$) on the image plane,
-$\mu_{2D}=\frac{KR_{k}^{\intercal}(\mu-T_{k})}{d},\ r_{2D}=\frac{fr}{d},\ d=e_{3}^{\intercal}R_{k}^{\intercal}(\mu-T_{k}),$
+[…]
 where $K$ is the intrinsic matrix, $R_{k}$ and $T_{k}$ are the rotational and translational components of the camera pose in the world frame at time $k$, $f$ is the focal length, and $e_{3}=[0,0,1]^{\intercal}$.
-For a given pixel $p$ in the image, its color $\mathcal{C(\cdot)}$ and depth $\mathcal{D(\cdot)}$ can be obtained as
-$\mathcal{C}(p)=\sum_{i=1}^{n}c_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p))\$
-and
-$\mathcal{D}(p)=\sum_{i=1}^{n}d_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p)),$
-where
-$f_{i}(p)=\alpha_{i}\exp\left\{-\frac{||p-\mu_{2D,i}||^{2}}{2r_{2D,i}^{2}}\right\}.$
-At every iteration, we render color image $\hat{y}_{c}$ and depth image $\hat{y}_{d}$. We set the function $\mathcal{L}$ (from eq. (2)) to
-$\mathcal{L}$ $=\frac{1}{N_{p}}\sum_{p\in y}\left(|\hat{y}_{d}-y_{d}|+\lambda_{1}|\hat{y}_{c}-y_{c}|\right)$ (4)
-$+\lambda_{2}(1-SSIM(\hat{y}_{c},y_{c}))$
+For a given pixel $p$ in the image, its color $\mathcal{C(\cdot)}$ can be obtained as
+[…]
+Similarly, the depth $\mathcal{D(\cdot)}$ can be obtained by replacing $c_{i}$ with $d_{i}$.
+At every iteration, we render color image $\hat{y}_{c}$ and depth image $\hat{y}_{d}$.
+We set the function $\mathcal{L}$ (from eq.
+(2)) to
+$\mathcal{L}=\frac{1}{N_{p}}\sum_{p\in y}\left(|\hat{y}_{d}-y_{d}|+\lambda_{1}|\hat{y}_{c}-y_{c}|\right)+\lambda_{2}(1-SSIM(\hat{y}_{c},y_{c}))$
 which is a weighted combination of the L1 loss on depth and rendered pixel colors and the structured similarity index measure (SSIM) to update the parameters of the Gaussians, and $N_{p}$ is the number of pixels in the image.
 IV Method
-Our proposed framework comprises the mapping module and the planning module, as illustrated in Fig.
+Our proposed framework comprises a mapping module and a planning module, as illustrated in Fig.
 2.
 The mapping module (Sec.
-IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the uncertainty of Gaussians in the map.
+IV-A) accumulates measurements and poses to reconstruct the environment and computes the uncertainty of Gaussians in the map.
 This is then passed to the planning module for planning guidance paths (Sec.
@@
 IV-A Mapping & Uncertainty Estimation
-We build the 3DGS mapping module upon the [14] framework.
-Similar to [14], we set all Gaussians to be isotropic.
+We build the 3DGS mapping module upon SplaTAM [8] with isotropic Gaussians.
 As in [8], we set $\lambda_{1}=0.4,\lambda_{2}=0.1$.
@@
 As in [2], we prune Gaussians with low opacity or large radius.
-Uncertainty Estimation. The objective of finding the next-best-view can then be expressed as
-$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathcal{I}(y_{k+1};\hat{G}|x_{k+1}).$ (5)
-The information gain $\mathcal{I}(\cdot)$ is defined as
-$\mathcal{I}$ $(y_{{k+1}};\hat{G}|x_{k+1})$ (6)
-$=\mathbb{H}(y_{k+1}|x_{k+1})-\mathbb{H}(y_{k+1}|\hat{G},x_{k+1})$
-where $\mathbb{H}(\cdot)$ is the entropy.
-We make the following assumptions in this work to approximate the information gain efficiently: (a) $\hat{G}\sim\mathcal{N}(\mu_{\hat{G}},\Sigma_{\hat{G}})$; (b) $y_{k+1}|\hat{G}\sim\mathcal{N}(h(x_{k+1},\hat{G}),\Sigma_{y})$; (c) constant isotropic variance $\Sigma_{y}$.
-From (a) and (b), the information gain can be simplified as
-$\mathcal{I}(y_{{k+1}};\hat{G}|x_{k+1})$ $=\frac{1}{2}\log\det\left(\frac{J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}+\Sigma_{y}}{\Sigma_{y}}\right)$ (7)
-$=\frac{1}{2}\log\det(I+\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal})$
-where $J_{k+1}=\frac{\partial h(x_{k+1},\hat{G})}{\partial G}$. Using the first order approximation of $\log\det$ about the identity, we can define a proxy metric for scoring next best views as
-$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathrm{Tr}(\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}).$ (8)
+Uncertainty Estimation. The objective for finding the next-best-view can then be expressed as maximizing the mutual information
+[…]
+[…]
+where ${H}(\cdot)$ is the Shannon entropy. We have
+$H(y_{k+1}\mid x_{k+1})=-\int p(y_{k+1}\mid x_{k+1})\log p(y_{k+1}\mid x_{k+1}))dy_{k+1}$
+with $p(y_{k+1}\mid x_{k+1})=\int p(y_{k+1}\mid\hat{G},x_{k+1})p(\hat{G})d\hat{G}$ and
+$H(y_{k+1}\mid$
+[…]
+$\biggl[\int p(y_{k+1}\mid\hat{G},x_{k+1})\log p(y_{k+1}\mid\hat{G},x_{k+1})\,dy_{k+1}\biggr]d\hat{G}$
+We make the following assumptions in this work to approximate the information gain efficiently: (a) $\hat{G}\sim\mathcal{N}(\mathrm{E}[G],\mathrm{Var}[G])$, where $G$ is the random variable that denotes the map; (b) $p(y_{k+1}\mid\hat{G})=\mathcal{N}(h(x_{k+1},\hat{G}),\Sigma_{y})$; with (c) constant isotropic variance $\Sigma_{y}$.
+For the marginal entropy ${H}(y_{k+1}\mid x_{k+1})$, given assumptions (a) and (b), and linearization of $h(\cdot)$ around $\mathrm{E}[G]$, we can approximate the distribution $p(y\mid x)=\mathcal{N}(h(x,\mathrm{E}[G]),J_{k+1}\mathrm{Var}[G]J_{k+1}^{\intercal})$, where $J_{k+1}=\partial h(x_{k+1},\hat{G})/\partial G$.
+Under assumptions (b) and (c), the conditional entropy does not depend on $G$.
+Recall the entropy of a $d_{Z}$ dimensional Gaussian distribution $\mathcal{N}(\mu_{Z},\Sigma_{Z})$ is $\frac{d_{Z}}{2}\log(2\pi e)+\frac{1}{2}\log\det(\Sigma_{Z})$.
+Dropping the constant terms, the information gain can be simplified to
+[…]
+Using the first order approximation of $\log\det$ about the identity, we can define a proxy metric for scoring next best view as
+[…]
 As in [24], assumption (c) allows us to simplify the optimization by ignoring the effect of $\Sigma^{-1}_{y}$.
@@
 Intuitively, this captures the set of Gaussians within the field of view at pose $x_{k+1}$.
-The key insight behind this approximation is that we estimate information gain using the current uncertainty, based on the assumption that local measurements lead to precise local map estimates.
+The key insight behind these approximations is that we estimate information gain using the current uncertainty, based on the assumption that local measurements lead to precise local map estimates.
 We show empirically that the exploration performance based on our metric (in Sec.
-V-A) is similar to the performance based on eq.
+V-A2) is similar to the performance based on eq.
 (7).
-We propose a heuristic for estimating uncertainty $\Sigma_{\hat{G}}$ based on the magnitude of the change in means ($\mu$) of the Gaussians, associating larger displacements with higher uncertainty.
-The motivation behind this lies in the fact that $\Sigma_{\hat{G}}$ can be approximated with the empirical Fisher Information matrix [43], $\Sigma_{\hat{G}}\approx\nabla_{G}\mathcal{L}(x_{k},y_{k})\nabla_{G}\mathcal{L}(x_{k},y_{k})^{\intercal}$, which in turn can be directly linked to the square of the updates of Gaussian parameters tuned by the gradient descent algorithm.
+We propose a heuristic for estimating uncertainty $\mathrm{Var}[G]$ based on the magnitude of the change in means ($\mu$) of the Gaussians, associating larger displacements with higher uncertainty.
+The motivation behind this is that $\mathrm{Var}[G]$ can be approximated with the empirical Fisher Information matrix [28], $\mathrm{Var}[G]\approx\nabla_{G}\mathcal{L}(x_{k},y_{k})\nabla_{G}\mathcal{L}(x_{k},y_{k})^{\intercal}$, which in turn can be directly linked to the square of the updates of Gaussian parameters tuned by the gradient descent algorithm.
 In practice, we observe that the square of the magnitude of the updates outlined above is sensitive to measurement noise (in Sec.
-V-A), so we use the magnitude of the change instead.
+V-A3), so we use the magnitude of the change instead.
 Empirically, we observe that the means of Gaussians located near boundaries between observed and unobserved space exhibit significant changes over successive updates, indicating high uncertainty, as visualized in Fig.
@@
 Consequently, directing exploration toward these high-uncertainty Gaussians is analogous to following frontiers in grid-based exploration, making this heuristic effective for autonomous exploration.
-Importantly, through careful bookkeeping, Gaussians that remain unchanged in the most recent update retain their previously computed displacement values, ensuring that our heuristic provides a consistent measure of each Gaussian’s movement over time.
+Through careful bookkeeping, Gaussians that remain unchanged in the most recent update retain their previously computed displacement values, ensuring that our heuristic provides a consistent measure of each Gaussian’s movement over time.
 In addition to the boundaries between observed and unobserved space, areas within observed space that have insufficient measurements also exhibit relatively high uncertainty.
-This further directs the robot to explore these areas and gather more information, thereby improving the overall map quality.
+This further directs the robot to explore these areas and gather more information, improving the overall map quality.
 IV-B Hierarchical Planning
@@
 We evenly partition the a priori known enclosing space of the desired map into cuboidal regions.
-At time ${k}$, for a region $o$, denote the cardinality of the Gaussians in the region by $M_{k}$.
-We compute the mean uncertainty of that region $\Omega_{k}=1/{M_{k}}\left(\sum_{\mu\in o}\|\mu_{k}-\mu_{{k-1}}\|_{2}\right).$ We formulate a high-level guidance path that allows us to (i) navigate to regions of high uncertainty; and (ii) utilize the known traversable space to plan long-range trajectories without computationally expensive collision checks.
+At time k{k}, for a region $o$, denote the cardinality of the Gaussians in the region by $M_{k}$.
+We compute the mean uncertainty of that region $\Omega_{k}=1/{M_{k}}\left(\sum_{\mu\in o}\|\mu_{k}-\mu_{{k-1}}\|_{2}\right).$
+We formulate a high-level guidance path that allows us to (i) navigate to regions of high uncertainty; and (ii) utilize the known traversable space to plan long-range trajectories without computationally expensive collision checks.
 At the high level of our planner, we construct a tree by incrementally adding nodes along the traveled path.
 The tree consists of two types of nodes: odometry nodes and viewpoint nodes.
+As the robot moves, it verifies the presence of a nearby odometry node; if none exist, a new odometry node is added to the tree and connected to the nearest existing odometry nodes.
 At each planning iteration, we sample a fixed number of viewpoints around the identified regions.
@@
 Trajectories are evaluated based on the sum of the utilities of the viewpoints in each trajectory given by
-$\xi(x,\hat{G})=\sum_{g\in\hat{G}_{H}}v(x,g)-\lambda_{\xi}\sum_{g\in\hat{G}_{L}}v(x,g).$ (9)
+[…]
 The first term in eq.
@@
 The second term penalizes observation of stable Gaussians in explored areas and encourages exploration of unseen parts of the environment.
-The two terms are weighted by the $\lambda_{\xi}$.
+The two terms are weighted by $\lambda_{\xi}$.
 Given the path generated by the high-level planner, we select the furthest point $p_{goal}$ along the path that lies within a local planning range $R_{H}$, and set it as the center of the goal region $\mathcal{X}_{goal}$.
@@
 Problem 1.
-Trajectory Candidate Generation. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
-[…]
+Trajectory Candidate Generation. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $0,\tau]$ that solve:
+$\min_{u(\cdot),\tau}\ \lambda_{t}\tau+\int_{0}^{\tau}u(t)^{T}u(t)dt$ (9)
 $s.t.\ \forall t\in[0,\tau],\ x(0)=x_{0},\ x(\tau)\in\mathcal{X}_{goal},$
 $\dot{x}(t)=f(x(t),u(t)),\ x(t)\in\mathcal{X}_{free},$
-$||v(t)||_{2}\leq v_{max},\ |\omega(t)|\leq\omega_{max}$
+[…]
 where $\lambda_{t}$ weights the time cost with the control effort, $v_{max}$ and $\omega_{max}$ are actuation constraints, and $x_{0}$ is the initial state.
@@
 Motion primitives are generated with fixed control inputs over a time interval with known initial states.
-Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $[0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs.
+Given the actuation constraints $v_{max}$ and $\omega_{max}$ of the robot, we uniformly generate $N_{v}\times N_{\omega}$ samples from $0,v_{max}]\times[-\omega_{max},\omega_{max}]$ as the finite set of control inputs.
 Subsequently, motion primitives are constructed given the dynamics model, the controls $u$, and time discretization.
@@
 This allows the real-time expansion of the search tree.
-Tree Search.
-The cost of each valid motion primitive is defined by Prob.
-1.
-Since the maximum velocity of the robot is bounded by $v_{max}$, we consider the minimum time heuristic as $h(p):=||p_{goal}-p||_{2}/v_{max}$.
-We use A* to search through the motion primitives tree and keep top $N_{traj}$ candidate trajectories for the information gain maximization.
-Information Gain Maximization.
+[…]
+Information Maximization.
 A state sequence $\{x_{i}({l\tau/L})\ |\ l\in\{0,\ldots,L\}\}$ containing the end state of each of $L$ segments in the $i$-th trajectory is used for evaluating the information along the trajectory.
@@
 The parameters of the mapper were set to the default configuration of [8] for the TUM dataset.
-For simulation experiments, we used 50 mapping iterations.
+For viewpoint selection experiments, we used 50 mapping iterations.
 For real-world experiments, to enable real-time mapping and collision avoidance, we reduced the number of mapping iterations to 10 and pruned Gaussians once per mapping sequence.
@@
 V Results
-V-A Simulation Experiments & Ablations
-Table I: Benchmark experiments on iTHOR test scenes with 100 steps
+V-A Information Metric Evaluation & Ablation
+V-A1 Experiment setup & Baselines
+To evaluate our proposed heuristic on viewpoint selection, we conducted simulation experiments with the AI2-THOR simulator [31] on the iTHOR dataset.
+In these experiments, time budget is not imposed, and we run experiments on all test scenes with 100 steps each.
+The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms.
+At each step, the agent is allowed to translate 0.5m in both $x$ and $y$ positions and a discrete set of possible orientations.
+These candidate positions, along with their corresponding yaw angles, define the full set of potential viewpoints.
+Each viewpoint is evaluated using an uncertainty metric, and the one with the highest uncertainty is selected as the next action.
+The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
+Table I: Information metric benchmark experiments on iTHOR
 Methods PSNR SSIM LPIPS RMSE t/step
@@
 RTGuIDE 22.946 0.820 0.305 0.343 0.013
-Table II: Ablation experiments on iTHOR test scenes with 100 steps
+Viewpoint selected using information metric from FisherRF.
+Table II: Information metric ablation experiments on iTHOR
 Methods PSNR SSIM LPIPS RMSE t/step
@@
 RTGuIDE-sq w. noise 21.445 0.779 0.351 0.386 0.013
-V-A1 Experiment setup & Baselines
-To evaluate our proposed heuristic, we conduct simulation experiments with the AI2-THOR simulator [46] on the iTHOR dataset.
-The iTHOR dataset consists of indoor scenes including kitchens, bedrooms, bathrooms and living rooms.
-Each step in the simulator involves a translation and rotation.
-The agent is allowed to translate 0.5m in both $x$ and $y$ positions.
-At each allowed position, we sample a set of candidate yaw angles to form a discrete set of possible orientations.
-These candidate positions, along with their corresponding yaw angles, define the full set of potential viewpoints.
-Each viewpoint is evaluated using an uncertainty metric, and the one with the highest uncertainty is selected as the next action.
-We ran experiments on all test scenes from iTHOR with 100 steps each.
-The simulations were run on a desktop computer with an AMD Ryzen Threadripper PRO 5975WX and NVIDIA RTX A4000 (16GB).
 For each scene, we synthesized novel views from the generated maps from a set of uniformly sampled test poses and computed the image metrics against the ground truth images.
@@
 Our heuristic performs comparably with FisherRF and outperforms Ensemble across all metrics, while being more than an order of magnitude faster in computation time.
-Since we are computing the uncertainty directly on the Gaussians instead of the image, we avoid the computation cost of rendering every sampled viewpoint.
-This efficiency in computation is crucial in a sampling-based planner where we evaluate potentially hundreds of viewpoints in each planning iteration.
+FisherRF renders images at every sampled viewpoint to approximate the information gain with the full map parameters, while we avoid such computational cost by computing the uncertainty directly on the parameters of the Gaussians.
+This trades off the approximation accuracy with efficiency.
+This efficiency in computation is crucial for real-time operation where we evaluate potentially hundreds of viewpoints in each planning iteration.
 V-A3 Ablation experiments
@@
 V-B Planner Performance
-We evaluate the necessity of our proposed GPU-based planning approach in enabling real-time planning.
-In particular, we conducted experiments on trajectory planning with a 5-meter horizon, evaluating both GPU-accelerated collision checking and a fully CPU-based implementation.
-As illustrated in Fig.
-3, the planning time of CPU-based planner increases significantly with the number of Gaussians.
-For example, a 40m $\times$ 40m outdoor parking lot can contain around $3\times 10^{6}$ Gaussians.
-In this case, the CPU-based planner requires 11.65 seconds, whereas the GPU-based planner takes 0.62 seconds, achieving 18$\times$ speedup.
-The GPU-based collision testing enables parallel growth of the search tree at each layer, along with simultaneous and efficient collision checks between all test points and Gaussians directly on the GPU.
 […]
 Figure 3: Time spent to plan a trajectory with a 5m horizon (including 129 collision checking points) versus the number of Gaussians in the map.
-V-C Real-world Experiments & Benchmarks
-Table III: Quantitative Results of Real-world Experiments.
-Methods Env. Budget (Time) [s] PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ mIoUa $\uparrow$
-Ensemble Indoor 1 72m2 300 6.27 0.204 0.839 2.051 0.071
-FisherRF 16.80 0.665 0.393 0.242 0.314
-RTGuIDE 17.83 0.737 0.334 0.202 0.338
 […]
-Ensemble Indoor 2 189m2 600 10.41 0.269 0.787 2.410 0.156
-FisherRF 14.88 0.588 0.451 1.170 0.248
-RTGuIDE 16.40 0.711 0.299 0.689 0.391
-[…]
-Ensemble Outdoor 1 252m2 600 10.59 0.392 0.637 1.705 0.549
-FisherRF 17.58 0.687 0.349 0.519 0.600
-RTGuIDE 20.00 0.744 0.296 0.369 0.622
-[…]
-Ensemble Outdoor 2 480m2 900 16.79 0.538 0.576 1.422 0.159
-FisherRF 19.44 0.680 0.439 0.773 0.177
-RTGuIDE 20.27 0.750 0.345 0.399 0.174
-[…]
+V-C Closed-loop Simulation Experiments
+Table III: Closed-loop Experiments on MP3D in Unity Simulator
+Methods PSNR SSIM LPIPS RMSE Coverage
+[dB] $\uparrow$ $\uparrow$ $\downarrow$ [m] $\downarrow$ [%] $\uparrow$
+Ensemble 10.71 0.24 0.76 2.14 38.93
+FisherRF 15.37 0.49 0.59 1.06 66.40
+RTGuIDE 16.32 0.56 0.54 0.71 92.36
+GT 18.13 0.65 0.47 0.50
+Table IV: Quantitative Results of Real-world Experiments.
+Methods Env. Budget (Time) [s] PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE [m] $\downarrow$ Coverage [%] $\uparrow$ mIoUa $\uparrow$
+Ensemble
+Indoor 1
+72m2
+300 6.27 0.204 0.839 2.051 98.48 0.071
+FisherRF 16.80 0.665 0.393 0.242 99.86 0.314
+RTGuIDE 17.83 0.737 0.334 0.202 99.95 0.338
+GT 20.56 0.805 0.237 0.158 0.420
+Ensemble
+Indoor 2
+189m2
+600 10.41 0.269 0.787 2.410 84.10 0.156
+FisherRF 14.88 0.588 0.451 1.170 72.80 0.248
+RTGuIDE 16.40 0.711 0.299 0.689 100.00 0.391
+GT 18.39 0.803 0.243 0.454 0.422
+Ensemble
+Outdoor 1
+252m2
+600 10.59 0.392 0.637 1.705 48.30 0.549
+FisherRF 17.58 0.687 0.349 0.519 85.81 0.600
+RTGuIDE 20.00 0.744 0.296 0.369 100.00 0.622
+GT 22.95 0.828 0.241 0.195 0.631
+Ensemble
+Outdoor 2
+480m2
+900 16.79 0.538 0.576 1.422 49.59 0.159
+FisherRF 19.44 0.680 0.439 0.773 60.95 0.177
+RTGuIDE 20.27 0.750 0.345 0.399 98.35 0.174
+GT 24.19 0.835 0.285 0.227 0.246
 a Indoor 1 classes: [floor, chair, table, refrigerator, cabinet, backpack, plants].
@@
 V-C1 Experiment setup & Baselines
-To evaluate the effectiveness of our proposed framework, we compared it with two baselines onboard the robot.
-We first implemented FisherRF [36] by constructing both a voxel map and a Gaussian splatting map online.
+We conducted experiments in a Unity-based simulator with ROS and a simulated Clearpath Jackal robot with an RGBD sensor and ground truth odometry.
+This set of experiments was designed to evaluate all methods with real-time and closed-loop operation, considering robot dynamics and collision avoidance.
+The original implementation of FisherRF [24] does not model dynamics or include real-world experimental results, so we implement a closed-loop version of it.
+We first construct both a voxel map and a Gaussian splatting map online.
 We detect and cluster frontiers from the voxel map and use A* to generate reference paths to frontier clusters.
 Camera views along each path are evaluated to compute information gain.
-To generate collision-free and dynamically feasible trajectories, we generate trajectories along waypoints in the path through MPL [45] in the voxel map considering the unicycle model for robot dynamics.
-We also implemented an Ensemble method and deployed it on the robot, similar to the framework of FisherRF.
-In the ensemble approach, uncertainty is computed as described in Sec.
+To generate collision-free and dynamically-feasible trajectories, we generate trajectories along waypoints in the path through MPL [30] in the voxel map.
+We similarly implement an Ensemble method as described in Sec.
 V-A1.
+We obtained the ground truth (GT) by tele-operating the robot to uniformly survey the environment.
+The collected data was used to build a 3DGS map for GT and also served as test poses for the other methods.
+This represents the upper bound of the evaluation metrics with the Gaussian splatting method.
+The exploration budget for all methods is set to approximately 5 times the duration needed to fully survey the environment with tele-operation.
+For the baseline methods, we set the voxel resolution to 5cm.
+We rendered 500 novel views at the ground truth test poses for evaluation.
+The same image metrics on the rendered RGB and depth images are used as in Sec.
+V-A.
+We evaluated all methods on 5 MP3D [34] scenes [17DRP5sb8fy, HxpKQynjfin, 2t7WUuJeko7, 8194nk5LbLH, YVUC4YcDtcY] with a time budget of 1500s, 480s, 1500s, 1200s and 1800s respectively.
+These scenes consist of multi-room indoor environments with varying size and clutter.
+V-C2 Results
+The averaged statistics of the experiments are presented in Table III.
+We note that there are signifcant errors in the meshes in MP3D which affect the reconstruction results for all methods including the ground truth.
+Our proposed method achieves 0.95dB and 5.61dB higher PSNR than FisherRF and Ensemble.
+The significantly higher coverage demonstrates the importance of efficient viewpoint selection in real-time missions with limited operational budget.
+V-D Real-world Experiments & Benchmarks
+V-D1 Experiment setup & Baselines
+To evaluate the effectiveness of our proposed framework in real-world settings, we compared it with two baselines onboard the robot.
+We followed the same setup as in the closed-loop simulation experiments for the baselines and the ground truth GT.
 We deployed our method and the baselines in four different real-world environments on a Clearpath Jackal robot outfitted with an AMD Ryzen 5 3600 and RTX 4000 Ada SFF.
@@
 We used [35] to provide odometry and truncated the depth measurements at 5 meters for all methods in all of our experiments.
-For the baseline methods, we set the voxel resolution to 5cm.
-To generate the ground truth data for evaluation, we teleoperated the robot to uniformly sample the environment.
-The exploration budget for all methods is set to $\sim$5x the time needed to fully survey the environment with teleoperation.
-We rendered 500 novel views according to the ground truth poses and compared them against the corresponding images.
-We computed the same image metrics on the rendered RGB and depth images as in the simulation experiments.
+The same evaluation setup from Sec.
+V-C is employed.
 To verify the usefulness of the generated Gaussian map representation for downstream tasks in robotics, we also performed an evaluation on the task of semantic segmentation.
 We used Grounded SAM 2 [36] to obtain segmentation masks of the rendered and groundtruth images and computed the mean Intersection over Union (mIoU).
-V-C2 Qualitative results
+V-D2 Qualitative results
 Figure 4: Visualization of onboard constructed map
@@
 Constructed in real time onboard the robot, our Gaussian splatting map provides a detailed representation of the environment.
-Being able to render photorealistic images from the map benefits downstream tasks such as semantic segmentation, which are analyzed in detail in the next section.
-V-C3 Quantitative results
+Being able to render photorealistic images from the map benefits downstream tasks such as semantic segmentation, which are analyzed in the next section.
+V-D3 Quantitative results
 As shown in Tab.
-III, in real-world scenarios when the operational budget is limited, our proposed framework constructs maps of higher quality than the baselines onboard the robot.
+IV, in real-world scenarios when the operational budget is limited, our proposed framework constructs maps of higher quality than the baselines onboard the robot.
 To validate that our approach generates maps of reasonable quality, we also present maps generated with the ground truth samples.
 Results on novel view image rendering show that the maps constructed with our approach achieve 0.8-2.4 higher PSNR than FisherRF and 3.5-11.6 higher PSNR than Ensemble.
-Furthermore, when evaluated on depth image recovery, our method reduces the RMSE of the rendered depth by 16.5%-48% compared to FisherRF, and by 71.4%-90.2% compared to Ensemble.
+Furthermore, when evaluated on depth reconstruction, our method reduces the RMSE of the rendered depth by 16.5%-48% compared to FisherRF, and by 71.4%-90.2% compared to Ensemble.
 These results reflect that our method outperforms the baselines in achieving good coverage of the environments while also optimizing for map quality during exploration.
 For semantic segmentation, our approach achieves better mIoU scores for most experiments, indicating higher fidelity of the rendered images compared to the baselines.
-V-C4 Discussion
+V-D4 Discussion
 We attribute the performance of our approach to two key factors.
@@
 In this work, we present a framework for real-time active exploration and mapping with Gaussian splatting.
-In future work, we wish to explore the use of Gaussian opacities to consider occluded Gaussians and improve the uncertainty estimates.
-Another future direction is to consider semantic features together with the Gaussian maps to perform complex tasks in the environment like object search and represent dynamic scenes in our framework.
+A future direction is to consider semantic features together with the Gaussian splatting maps to perform complex tasks in the environment like object search and represent dynamic scenes in our framework.
 In this work, we study the active mapping problem which assumes perfect state estimation.